### 05 Jan 2024 v1

# Nonuniform Sobolev Spaces

Abstract.
We study nonuniform Sobolev spaces, i.e., spaces of functions whose partial derivatives lie in possibly different Lebesgue spaces. Although standard proofs do not apply, we show that nonuniform Sobolev spaces share similar properties as the classical ones. These spaces arise naturally in the study of certain PDEs. For instance, we illustrate that nonuniform fractional Sobolev spaces are useful in the study of convergence of Schrödinger operators. In this work we extend recent advances on the convergence of Schrödinger operators to nonuniform fractional Sobolev spaces.
Key words and phrases:
Sobolev spaces, Nonuniform Sobolev spaces, fractinoal Sobolev spaces, embedding theorems, Schrödinger operators.
2020 Mathematics Subject Classification:
46E35
T. Chen was supported by the National Natural Science Foundation of China (No. 12271267) and the Fundamental Research Funds for the Central Universities.
L. Grafakos was supported by a Simons Grant (No. 624733).
W. Sun was supported by the National Natural Science Foundation of China (No. 12171250 and U21A20426).
1. Introduction
The theory of Sobolev spaces plays an import role in the study of partial differential equations and many other fields. Given a positive integer $m$, a positive number $p\in[1,\infty]$ and an open set $\Omega\subset\mathbb{R}^{N}$, we denote by
$W^{m,p}(\Omega)=\{f\in L^{p}(\Omega):\,D^{\alpha}f\in L^{p}(\Omega),|\alpha|\leq m\},$
the classical Sobolev space, where $\alpha=(\alpha_{1},\ldots,\alpha_{N})$ is a multi-index, $|\alpha|=\sum_{i=1}^{N}\alpha_{i}$, and $D^{\alpha}$ denotes the weak partial derivatives, that is, partial derivatives in the sense of distributions.
Since a function and its derivatives might have different properties, it is not necessary for them to lie in the same Lebesgue space. In this paper, we consider nonuniform Sobolev spaces on $\mathbb{R}^{N}$, i.e., spaces for which a function and its derivatives belong to different Lebesgue spaces.
Definition 1.1.
Let $\Omega\subset\mathbb{R}^{N}$ be an open set. For $k\geq 1$ and $\vec{p}=(p_{0},\ldots,p_{k})\in[1,\infty]^{k+1}$, the nonuniform Sobolev space $W_{k}^{\vec{p}}(\Omega)$ consists of all measurable functions $f$ for which $\partial^{\alpha}f\in L^{p_{|\alpha|}}(\Omega)$, where $|\alpha|\leq k$. For $f\in W_{k}^{\vec{p}}$, define its norm by
$\|f\|_{W_{k}^{\vec{p}}}=\sum_{|\alpha|\leq k}\|D^{\alpha}f\|_{L^{p_{|\alpha|}}}.$
These spaces naturally arise in the study of certain PDEs. In particular, some equations have solutions in nonuniform Sobolev spaces but have no solution in the classical Sobolev spaces. For example, consider positive solutions of the critical $p$-Laplace equation
$\Delta_{p}u+u^{p^{*}-1}=0$ (1.1)
in $\mathbb{R}^{N}$, where $1<p<N$, $p^{*}=Np/(N-p)$ and $\Delta_{p}u=\mathrm{div}(|\nabla u|^{p-2}\nabla u)$.
This equation is well studied in the literature. It was shown by Damascelli, Merchán, Montoro and Sciunzi [15], Sciunzi [29] and Vétois [34] that if a function $u$ in the class
$\mathcal{D}^{1,p}(\mathbb{R}^{N})=\{u\in L^{p^{*}}(\mathbb{R}^{N}):\,\nabla u\in L^{p}(\mathbb{R}^{N})\}$
is a solution of (1.1), then it is of the form
$u(x)=U_{\lambda,x_{0}}(x)=\left(\frac{\lambda^{1/(p-1)}N^{1/p}((N-p)/(p-1))^{{(p-1)}/p}}{\lambda^{p/(p-1)}+|x-x_{0}|^{p/(p-1)}}\right)^{(N-p)/p},$
where $\lambda$ is a positive constant and $x_{0}$ is a point in $\mathbb{R}^{N}$. Moreover, Catino, Monticelli and Roncoroni [10] provided the classification of positive solutions to the critical $p$-Laplace equation. See also Ciraolo, Figalli and Roncoroni [12] for solutions in convex cones.
Observe that $\mathcal{D}^{1,p}$ itself is the nonuniform Sobolev space $W_{1}^{(p^{*},p)}$. Since $|U_{\lambda,x_{0}}(x)|\approx|x|^{-(N-p)/(p-1)}$ whenever $|x|$ is large enough, it is easy to see that (1.1) has a positive solution in $W^{1,p}$ if and only if $1<p<N^{1/2}$. However, we can increase the integrability index $p$ of the solution if we consider nonuniform Sobolev spaces. Indeed, we obtain that (1.1) has a positive solution in $W_{1}^{(p_{0},p)}$ whenever $1<p<N$ and $p_{0}>N(p-1)/(N-p)$.
On the other hand, for $1\leq p_{1}<N$ and $f$ in the classical Sobolev space $W^{1,p_{1}}$, the Sobolev inequality says that
$\|f\|_{L^{Np_{1}/(N-p_{1})}}\leq C_{N,p_{1}}\|\nabla f\|_{L^{p_{1}}}.$ (1.2)
This is proved by showing that
$\|f\|_{L^{Np_{1}/(N-p_{1})}}^{(N-1)p_{1}/(N-p_{1})}\leq C_{N,p_{1}}\|\nabla f\|_{L^{p_{1}}}\|f\|_{L^{Np_{1}/(N-p_{1})}}^{N(p_{1}-1)/(N-p_{1})}$ (1.3)
for compactly supported differentiable functions $f$. We refer to [16] for details.
For the nonuniform case, that is, for $f\in C^{1}\cap W_{k}^{\vec{p}}$ with $\vec{p}=(p_{0},p_{1})$, both (1.2) and (1.3) are still true. However, their proofs are quite different from the classical uniform case. In the classical case, one first obtains the density of compactly supported differentiable functions, then derives (1.2) from (1.3) for such functions, and finally extends (1.2) to all functions in $W^{1,p_{1}}$ by density. For the nonuniform case, compactly supported differentiable functions are still dense in $W_{1}^{\vec{p}}$. However, the embedding inequality is required in the proof. So we have to prove (1.2) directly for functions which is not compactly supported. In this case, (1.2) is not a straightforward consequence of (1.3): we have to show first that $f\in L^{Np_{1}/(N-p_{1})}$.
Another application of nonuniform Sobolev spaces is to study the convergence of Schrödinger operators defined by
$e^{it(-\Delta)^{a/2}}f(x):=\frac{1}{(2\pi)^{N}}\int_{\mathbb{R}^{N}}e^{i(x\cdot\omega+t|\omega|^{a})}\hat{f}(\omega)\mathrm{d}\omega,$
where $a>1$ is a constant. It is well known that for $f$ nice enough, $e^{it(-\Delta)^{a/2}}f(x)$ is the solution of the fractional Schrödinger equation
$\begin{cases}i\partial_{t}u+(-\Delta_{x})^{a/2}u=0,&(x,t)\in\mathbb{R}^{N}\times R,\\ u(x,0)=f(x),&x\in\mathbb{R}^{N}.\end{cases}$
For the case $N=1$ and $a=2$, Carleson [9] studied the convergence of $e^{-it\Delta}f(x)$ as $t$ tends to 00 for functions $f$ in the Sobolev space
$H^{s}(\mathbb{R}^{N}):=\{f\in L^{2}:\,(1+|\omega|^{2})^{s/2}\hat{f}(\omega)\in L^{2}\}.$
It was shown [9] that when $a=2$,
$\lim_{t\rightarrow 0}e^{it(-\Delta)^{a/2}}f(x)=f(x),\quad a.e.$ (1.4)
for all $f\in H^{s}(\mathbb{R})$ if $s\geq 1/4$.
Since then, many works have been done on this topic. Dahlberg and Kenig [14] provided counterexamples indicating that the range $s\geq 1/4$ is sharp for $N=1$. And Sjölin [31] extended this result to the case $a>1$.
For $a=2$ and higher dimensions $N\geq 2$, Sjölin [31] and Vega [33] proved that (1.4) is true when $s>1/2$. For the case $N=2$, this result was improved to $s>3/8$ by Lee [23]. Bourgain  [3, 4] proved that $s>1/2-1/(4N)$ is sufficient and $s\geq 1/2-1/(2N+2)$ is necessary for the convergence. Du, Guth and Li [18] proved that $s>1/2-1/(2N+2)$ is sufficient for the dimension $N=2$ and Du and Zhang [19] showed that it is also true for general $N\geq 3$.
For the general case $a>1$ and $N\geq 2$, Sjölin [31] proved that (1.4) is valid when $s>1/2$. Prestini [26] showed that $s\geq 1/4$ is necessary and sufficient for radial functions $f$ and dimensions $N\geq 2$. Cho and Ko [11] proved that $s>1/3$ is sufficient for the dimension $N=2$.
Related works also include the non-tangential convergence by Shiraki [30], Yuan, Zhao and Zheng [36], Li, Wang and Yan [24, 25], the convergence along curves by Cao and Miao [8] and Zheng [37], and the convergence rate by Cao, Fan and Wang [7]; see also the works by Cowling [13], Walther [35], and Rogers and Villarroya [27] for the case $a<1$.
Although the range $s\geq 1/4$ of the index $s$ is sharp for $N=1$, and $s>1/2-1/(2N+2)$ is also sharp up to the endpoint for $N\geq 2$, we show that the result can be further extended when nonuniform Sobolev spaces are considered.
Suppose that $10<s<$ 1 and $1\leq p<\infty$. Recall that the classical fractional Sobolev space $W_{s}^{p}$ consists of all measurable functions $f$ for which
$\|f\|_{W_{s}^{p}}:=\|f\|_{L^{p}}+[f]_{W_{s}^{p}}<\infty,$
where
$[f]_{W_{s}^{p}}:=\bigg{(}\iint_{\mathbb{R}^{N}\times\mathbb{R}^{N}}\frac{|f(x)-f(y)|^{p}}{|x-y|^{N+sp}}\mathrm{d}x\mathrm{d}y\bigg{)}^{1/p}$
is the Gagliardo seminorm of $f$. Fractional Sobolev spaces were introduced by Aronszajn [2], Gagliardo [20] and Slobodeckiĭ [32]. We refer to [1, 5, 17] for an overview of fractional Sobolev spaces and [6, 22] for the limit of norms of fractional Sobolev spaces.
We now provide a formal definition of nonuniform fractional Sobolev spaces.
Definition 1.2.
Given $10<s<$ 1 and $\vec{p}=(p_{0},p_{1})\in[1,\infty)^{2}$, the nonuniform fractional Sobolev space is defined by
$W_{s}^{\vec{p}}(\mathbb{R}^{N})$ =$=\{f:\,\|f\|_{W_{s}^{\vec{p}}}:=\|f\|_{L^{p_{0}}}+[f]_{W_{s}^{p_{1}}}<\infty\}.$
For the case $s>1$, set $\nu_{s}=s-\lfloor s\rfloor$ and $\vec{p}=(p_{0},\ldots,p_{\lceil s\rceil})$. If $s$ is not an integer, define
$W_{s}^{\vec{p}}(\mathbb{R}^{N})=\Big{\{}f:\,\|f\|_{W_{s}^{\vec{p}}}:=\sum_{l=0}^{\lfloor s\rfloor}\sum_{|\alpha|=l}\|D^{\alpha}f\|_{L^{p_{l}}}+\sum_{|\alpha|=\lfloor s\rfloor}[D^{\alpha}f]_{W_{\nu_{s}}^{p_{\lceil s\rceil}}}<\infty\Big{\}}.$
Here $\lfloor s\rfloor$ stands for the greatest integer which is less than or equal to $s$, and $\lceil s\rceil$ stands for the ceiling, i.e., the least integer which is greater than or equal to $s$.
Recall that if $p_{0}=\ldots=p_{\lceil s\rceil}=2$, then $W_{s}^{(2,\ldots,2)}(\mathbb{R}^{N})$ coincides with $H^{s}(\mathbb{R}^{N})$. We refer to [16, Proposition 4.17] for a proof.
The introduction and study of these spaces is motivated by the fact that they provide stronger results than classical Sobolev spaces. As an application we show that if the convergence of Schrödinger operators holds for all functions in $H^{s}$ for some $s>0$, then it also holds for all functions in $W_{s}^{\vec{p}}$ with the same index $s$ if $p_{\lceil s\rceil}=2$. Moreover, we also obtain convergence results in the case $21<p_{\lceil s\rceil}<$ 2.
The paper is organized as follows. In Section 2, we show that compactly supported infinitely differentiable functions are dense in nonuniform Sobolev spaces. We also obtain the Sobolev inequality for functions in nonuniform Sobolev spaces. In Section 3, we focus on nonuniform fractional Sobolev spaces and present an embedding theorem for such spaces. And in Section 4, we give an application of nonuniform Sobolev spaces. We prove the almost everywhere convergence of Schrödinger operators for a large class of functions, extending known results.
[…]
For a tempered distribution $f$ and a function $\varphi$ in the Schwartz class $\mathscr{S}$, the notation $\langle f,\varphi\rangle$ stands for the value of the action of $f$ on $\varphi$.
2. Embedding Inequalities for Nonuniform Sobolev Spaces
Denote by $C_{c}^{\infty}(\mathbb{R}^{N})$ the function space consisting of all compactly supported infinitely differentiable functions. In this section we obtain the density of $C_{c}^{\infty}(\mathbb{R}^{N})$ in nonuniform Sobolev spaces. We then obtain a Sobolev embedding theorem for nonuniform Sobolev spaces; and the aforementioned density is a crucial tool in this embedding. Both of these results seem to require proofs that are quite different from the classical case. For the nonuniform case, we need first to prove the density for special indices. Then we deduce an embedding result, with which we finally obtain the density of $C_{c}^{\infty}(\mathbb{R}^{N})$ for full indices.
We begin with a simple lemma.
Lemma 2.1.
For any $k\geq 1$ and $\vec{p}\in[1,\infty)^{k+1}$, the space $C^{\infty}\cap W_{k}^{\vec{p}}(\mathbb{R}^{N})$ is dense in $W_{k}^{\vec{p}}(\mathbb{R}^{N})$.
Proof.
Take some $f\in W_{k}^{\vec{p}}$ and $\varphi\in C_{c}^{\infty}$ such that $\varphi$ is nonnegative and $\|\varphi\|_{L^{1}}=1$. Set
$g_{\lambda}(x)=\int_{\mathbb{R}^{N}}f(y)\frac{1}{\lambda^{N}}\varphi(\frac{x-y}{\lambda})\mathrm{d}y,\qquad\lambda>0.$
We have $g_{\lambda}\in C^{\infty}$. For any multi-index $\alpha$ with $|\alpha|\leq k$, we have
$\partial^{\alpha}g_{\lambda}(x)$ $=\int_{\mathbb{R}^{N}}\partial^{\alpha}f(y)\frac{1}{\lambda^{N}}\varphi(\frac{x-y}{\lambda})\mathrm{d}y$ (2.1)
$=\int_{\mathbb{R}^{N}}\partial^{\alpha}f(x-\lambda y)\varphi(y)\mathrm{d}y.$
It follows from Minkowski’s inequality and Lebesgue’s dominated convergence theorem that
$\lim_{\lambda\rightarrow 0}\|\partial^{\alpha}g_{\lambda}-\partial^{\alpha}f\|_{L^{p_{|\alpha|}}}\leq\lim_{\lambda\rightarrow 0}\int_{\mathbb{R}^{N}}\|\partial^{\alpha}f(\cdot-\lambda y)-\partial^{\alpha}f\|_{L^{p_{|\alpha|}}}\varphi(y)\mathrm{d}y=0.$
This completes the proof. ∎
The proof of the density of $C_{c}^{\infty}$ in nonuniform Sobolev spaces is split in two cases. First, we consider special indices.
Lemma 2.2.
Suppose that $k\geq 1$ and $\vec{p}=(p_{0},\ldots,p_{k})$ with $1/p_{i}\leq 1/p_{i-1}+1/N$, $1\leq i\leq k$. Then the space $C_{c}^{\infty}(\mathbb{R}^{N})$ is dense in $W_{k}^{\vec{p}}(\mathbb{R}^{N})$.
Proof.
By Lemma 2.1, for any $f\in W_{k}^{\vec{p}}$ and $\varepsilon>0$, there is some $g\in C^{\infty}\cap W_{k}^{\vec{p}}$ such that
$\|g-f\|_{W_{k}^{\vec{p}}}<\varepsilon.$
Moreover, if follows from (2.1) and Young’s inequality that we may choose the function $g$ such that
$\partial^{\gamma}g\in L^{r},\qquad\forall\,\,\,r\geq p_{|\gamma|}.$ (2.2)
The proof will be complete if we can show that there is some $\tilde{g}\in C_{c}^{\infty}$ such that
$\|g-\tilde{g}\|_{W_{k}^{\vec{p}}}<C^{\prime}\varepsilon.$ (2.3)
Take some $\psi\in C_{c}^{\infty}$ such that $\psi(x)=1$ whenever $|x|<1$. Set $g_{n}(x)=\psi(x/n)g(x)$.
It remains to show that the sequence $\{g_{n}:\,n\geq 1\}$ converges to $g$ in $W_{k}^{\vec{p}}$.
We see from the choice of $\psi$ that $\{g_{n}:\,n\geq 1\}$ converges to $g$ in $L^{p_{0}}$. On the other hand, fix some multi-index $\alpha$ with $1\leq|\alpha|\leq k$. Then $\partial^{\alpha}g_{n}(x)$ is the sum of $\psi(x/n)\partial^{\alpha}g(x)$ and terms like $(1/n^{|\beta|})\partial^{\beta}\psi(x/n)\partial^{\gamma}g(x)$, where $|\beta|+|\gamma|=|\alpha|$ and $|\beta|\geq 1$.
If $p_{|\alpha|}\geq p_{|\gamma|}$, we deduce from (2.2) that
$\Big{\|}\frac{1}{n^{|\beta|}}\partial^{\beta}\psi(\frac{\cdot}{n})\partial^{\gamma}g\Big{\|}_{L^{p_{|\alpha|}}}$ $\leq\frac{1}{n^{|\beta|}}\cdot\|\partial^{\beta}\psi\|_{L^{\infty}}\|\partial^{\gamma}g1_{\{|x|\geq n\}}\|_{L^{p_{|\alpha|}}}$
$\rightarrow 0,\qquad\mathrm{as}\,\,n\rightarrow\infty.$
If $p_{|\alpha|}<p_{|\gamma|}$, there is some $r>1$ such that
$\frac{1}{p_{|\alpha|}}=\frac{1}{p_{|\gamma|}}+\frac{1}{r}.$
Applying Hölder’s inequality, we have
$\Big{\|}\frac{1}{n^{|\beta|}}\partial^{\beta}\psi(\frac{\cdot}{n})\partial^{\gamma}g\Big{\|}_{L^{p_{|\alpha|}}}$ $\leq\frac{1}{n^{|\beta|}}\cdot\|\partial^{\beta}\psi(\frac{\cdot}{n})\|_{L^{r}}\|\partial^{\gamma}g1_{\{|x|\geq n\}}\|_{L^{p_{|\gamma|}}}$
$=\frac{1}{n^{|\beta|-N/r}}\cdot\|\partial^{\beta}\psi\|_{L^{r}}\|\partial^{\gamma}g1_{\{|x|\geq n\}}\|_{L^{p_{|\gamma|}}}.$ (2.4)
Since $1/p_{i}\leq 1/p_{i-1}+1/N$ for all $1\leq i\leq k$, we have
$\frac{1}{r}=\frac{1}{p_{|\alpha|}}-\frac{1}{p_{|\gamma|}}\leq\frac{|\alpha|-|\gamma|}{N}=\frac{|\beta|}{N}.$
Consequently, $|\beta|-N/r\geq 0$. It follows from (2.4) and (2.2) that
$\Big{\|}\frac{1}{n^{|\beta|}}\partial^{\beta}\psi(\frac{\cdot}{n})\partial^{\gamma}g\Big{\|}_{L^{p_{|\alpha|}}}\rightarrow 0,\qquad\mathrm{as}\,\,n\rightarrow\infty.$ (2.5)
Hence the sequence $\{g_{n}:\,n\geq 1\}$ converges to $g$ in $W_{k}^{\vec{p}}$ and (2.3) is valid. ∎
To prove the density of $C_{c}^{\infty}$ for full indices, we first establish an embedding theorem. For classical Sobolev spaces, the embedding inequality
$\|\varphi\|_{L^{N/(N-1)}}\leq C\|\nabla\varphi\|_{L^{1}}$
is first proved for functions in $C_{c}^{\infty}$, and then for general functions in $W^{1,1}$ by the density of $C_{c}^{\infty}$ in $W^{1,1}$. We refer to [16, eq. (2.38)] for details.
In the nonuniform case, we have no such density result at the moment. So we need to prove it directly for functions in $C^{\infty}\cap L^{p_{0}}(\mathbb{R}^{N})$ for some $p_{0}>0$.
Lemma 2.3.
Suppose that $N\geq 2$.
  1. (i)
    For any $\varphi\in C^{\infty}\cap L^{p_{0}}(\mathbb{R}^{N})$ with $0<p_{0}<\infty$,
    $\|\varphi\|_{L^{N/(N-1)}}\leq\frac{1}{N}\sum_{|\alpha|=1}\|\partial^{\alpha}\varphi\|_{L^{1}}.$ (2.6)
    Moreover, the above inequality is also true for any $f\in W_{1}^{\vec{p}}$ with $\vec{p}=(p_{0},1)$ and $1\leq p_{0}<\infty$.
  2. (ii)
    Suppose that $\vec{p}=(p_{0},p_{1})$, where $1\leq p_{0}<\infty$ and $1\leq p_{1}<N$. For any $f\in W_{1}^{\vec{p}}(\mathbb{R}^{N})$, we have
    $\|f\|_{L^{Np_{1}/(N-p_{1})}}\leq C_{N,p_{1}}\|\nabla f\|_{L^{p_{1}}}.$ (2.7)
    Hence for all $q$ between $p_{0}$ and $Np_{1}/(N-p_{1})$, we have
    $\|f\|_{L^{q}}\leq C\|f\|_{W_{1}^{\vec{p}}}.$ (2.8)
Proof.
(i)   The proof is similar to that for [16, eq. (2.38)]. Here we only provide a sketch with emphasis on the difference.
Fix some $\varphi\in C^{\infty}\cap L^{p_{0}}$. If the right-hand side of (2.6) equals infinity, then (2.6) is certainly valid. So we only need to consider the case $\|\partial^{\alpha}\varphi\|_{L^{1}}<\infty$ for all multi-indices $\alpha$ with $|\alpha|=1$.
For each index $i$, denote $\tilde{x}_{i}=(x_{1},\ldots,x_{i-1},x_{i+1},\ldots,x_{N})$. Since $\varphi\in L^{p_{0}}$, we have
$\int_{\mathbb{R}^{N-1}}\bigg{(}\int_{\mathbb{R}}|\varphi(x)|^{p_{0}}\mathrm{d}x_{i}\bigg{)}\,\mathrm{d}\tilde{x}_{i}=\|\varphi\|_{L^{p_{0}}}^{p_{0}}<\infty.$
Hence there are measurable sets $A_{i}\subset\mathbb{R}^{N-1}$ of measure zero such that
$\int_{\mathbb{R}}|\varphi(x)|^{p_{0}}\mathrm{d}x_{i}<\infty,\qquad\forall\,\,\tilde{x}_{i}\in\mathbb{R}^{N-1}\setminus A_{i}.$
Consequently, for each $\tilde{x}_{i}$ there is a sequence $\{a_{n}:\,n\geq 1\}\subset\mathbb{R}$ (depending on $\tilde{x}_{i}$) such that $\lim_{n\rightarrow\infty}a_{n}=-\infty$ and $\lim_{n\rightarrow\infty}\varphi(x+(a_{n}-x_{i})e_{i})=0$. Since for each $x_{i}$ we have
$\varphi(x)-\varphi(x+(a_{n}-x_{i})e_{i})=\int_{a_{n}}^{x_{i}}\partial_{x_{i}}\varphi(x+(t-x_{i})e_{i})\mathrm{d}t,$
by the Fundamental Theorem of Calculus, letting $n\rightarrow\infty$, we obtain
$\varphi(x)=\int_{-\infty}^{x_{i}}\partial_{x_{i}}\varphi(x+(t-x_{i})e_{i})\mathrm{d}t,\qquad\forall\,x_{i}\in\mathbb{R},\,\,\forall\,\tilde{x}_{i}\in\mathbb{R}^{N-1}\setminus A_{i}.$
Here and henceforth $e_{i}=(0,\ldots,1,\ldots,0)$ with $1$ only on the $i$th entry and 00 elsewhere. It follows that
$|\varphi(x)|\leq\int_{\mathbb{R}}\left|\partial_{x_{i}}\varphi(x+(t-x_{i})e_{i})\right|\mathrm{d}t,\qquad\forall\,x\in\mathbb{R}^{N}\setminus B_{i},$ (2.9)
where
$B_{i}=\{x\in\mathbb{R}^{N}:\,x_{i}\in\mathbb{R},\tilde{x}_{i}\in A_{i}\}$
is of measure zero in $\mathbb{R}^{N}$. Observe that the integral in (2.9) is independent of $x_{i}$.
For $1\leq i\leq N$, define the function $F_{i}$ on $\mathbb{R}^{N-1}$ by
$F_{i}(\tilde{x}_{i})=\int_{\mathbb{R}}\left|\partial_{x_{i}}\varphi(x+(t-x_{i})e_{i})\right|\mathrm{d}t.$
We have
$|\varphi(x)|^{N/(N-1)}\leq\prod_{i=1}^{N}F_{i}(\tilde{x}_{i})^{1/(N-1)},\qquad\forall x\in\mathbb{R}^{N}\setminus\bigcup_{i=1}^{N}B_{i}.$
Now following the same arguments as that in [16, Page 76] we obtain
$\|\varphi\|_{L^{N/(N-1)}}\leq\frac{1}{N}\sum_{|\alpha|=1}\|\partial^{\alpha}\varphi\|_{L^{1}}.$
For the general case, i.e., $f\in W_{1}^{\vec{p}}$ with $p_{1}=1$, since $C^{\infty}\cap W_{1}^{\vec{p}}$ is dense in $W_{1}^{\vec{p}}$, there is a sequence $\{\varphi_{n}:\,n\geq 1\}\subset C^{\infty}\cap W_{1}^{\vec{p}}$ which is convergent to $f$ in $W_{1}^{\vec{p}}$, as well as on $\mathbb{R}^{N}$ almost everywhere.
For each $n$, we have
$\|\varphi_{n}\|_{L^{N/(N-1)}}\leq\frac{1}{N}\sum_{|\alpha|=1}\|\partial^{\alpha}\varphi_{n}\|_{L^{1}}.$
Letting $n\rightarrow\infty$, we apply Fatou’s lemma to finally deduce
$\|f\|_{L^{N/(N-1)}}\leq\frac{1}{N}\sum_{|\alpha|=1}\|\partial^{\alpha}f\|_{L^{1}}.$
(ii)   We see from the embedding theorem for Sobolev’s spaces (see [16, eq. (2.46)]) that for any $f\in C_{c}^{\infty}(\mathbb{R}^{N})$,
$\|f\|_{L^{Np_{1}/(N-p_{1})}}\leq C\|\nabla f\|_{L^{p_{1}}}.$
If $1/p_{1}\leq 1/p_{0}+1/N$, then $C_{c}^{\infty}$ is dense in $W_{1}^{\vec{p}}$, thanks to Lemma 2.2. Similar arguments as the previous case we get that the above inequality is true for all $f\in W_{1}^{\vec{p}}$.
It remains to consider the case $1/p_{1}>1/p_{0}+1/N$.
Take $f\in W_{1}^{\vec{p}}$. Set $\vec{\tilde{p}}=(\tilde{p}_{0},p_{1})$, where $1/\tilde{p}_{0}=1/p_{1}-1/N$. It is easy to see that $1<\tilde{p}_{0}<p_{0}$. We conclude that $f\in W_{1}^{\vec{\tilde{p}}}$, for which we only need to show that $f\in L^{\tilde{p}_{0}}$.
To this end, set $r=p_{0}/p^{\prime}_{1}+1$ and $h=|f|^{r-1}f$. We have $\nabla h=r|f|^{r-1}\nabla f$. Since $f\in L^{p_{0}}$, $\nabla f\in L^{p_{1}}$ and $(r-1)/p_{0}+1/p_{1}=1$, we see from Hölder’s inequality that
$\|\nabla h\|_{1}\leq r\||f|^{r-1}\|_{L^{p_{0}/(r-1)}}\|\nabla f\|_{L^{p_{1}}}=r\|f\|_{L^{p_{0}}}^{r-1}\|\nabla f\|_{L^{p_{1}}}.$
It follows from (2.6) that
$\|h\|_{L^{N/(N-1)}}$ $\leq C\|\nabla h\|_{L^{1}}\leq Cr\|f\|_{L^{p_{0}}}^{r-1}\|\nabla f\|_{L^{p_{1}}}.$
Denote $q=rN/(N-1)$. Observe that $\|h\|_{L^{N/(N-1)}}=\|f\|_{L^{q}}^{r}$. We have
$\|f\|_{L^{q}}\leq\|f\|_{L^{p_{0}}}^{1-1/r}\Big{(}Cr\|\nabla f\|_{L^{p_{1}}}\Big{)}^{1/r}.$ (2.10)
Since $1/p_{1}>1/p_{0}+1/N$, we have
$\frac{q}{p_{0}}=\frac{N}{N-1}\Big{(}\frac{1}{p_{0}}+1-\frac{1}{p_{1}}\Big{)}<1.$
On the other hand, it follows from $1/p_{0}<1/p_{1}-1/N$ that
$\frac{1}{p^{\prime}_{1}}<\frac{p_{0}}{p^{\prime}_{1}}\Big{(}\frac{1}{p_{1}}-\frac{1}{N}\Big{)}.$
Hence,
$1-\frac{1}{N}<\frac{1}{p_{1}}-\frac{1}{N}+\frac{p_{0}}{p^{\prime}_{1}}\Big{(}\frac{1}{p_{1}}-\frac{1}{N}\Big{)}.$
Therefore we have
$\frac{1}{q}=\frac{N-1}{N}\cdot\frac{1}{1+p_{0}/p^{\prime}_{1}}<\frac{1}{p_{1}}-\frac{1}{N}.$
Set $q_{0}=p_{0}$. For $n\geq 1$, define $r_{n}$ and $q_{n}$ recursively by
$r_{n}=\frac{q_{n-1}}{p^{\prime}_{1}}+1\mbox{\quad and \quad}q_{n}=\frac{r_{n}N}{N-1}.$
We see from above arguments that both $\{q_{n}:\,n\geq 1\}$ and $\{r_{n}:\,n\geq 1\}$ are decreasing, $1/p_{1}>1/q_{n}+1/N$ and
$\|f\|_{L^{q_{n}}}\leq\|f\|_{L^{q_{n-1}}}^{1-1/r_{n}}\Big{(}Cr_{n}\|\nabla f\|_{L^{p_{1}}}\Big{)}^{1/r_{n}}.$ (2.11)
Since $q_{n}$ is bounded from below, the limits $\tilde{r}:=\lim_{n\rightarrow\infty}r_{n}$ and $\tilde{q}:=\lim_{n\rightarrow\infty}q_{n}$ exist. Moreover, $\tilde{q}=\tilde{r}N/(N-1)>\tilde{r}\geq 1$.
Take some constant $\varepsilon$ small enough such that $\varepsilon\cdot Cr_{1}\|\nabla f\|_{L^{p_{1}}}<1$. Since $r_{n}\leq r_{1}$, we have $\varepsilon\cdot Cr_{n}\|\nabla f\|_{L^{p_{1}}}<1$ for all $n\geq 1$. Set $\tilde{f}=\varepsilon f$ and substitute $\tilde{f}$ for $f$ in (2.11), we get
$\|\tilde{f}\|_{L^{q_{n}}}\leq\|\tilde{f}\|_{L^{q_{n-1}}}^{1-1/r_{n}},\qquad n\geq 1.$
Applying the above inequality recursively, we obtain
$\|\tilde{f}\|_{L^{q_{n}}}$ $\leq\|\tilde{f}\|_{L^{q_{n-2}}}^{(1-1/r_{n})(1-1/r_{n-1})}$
[…]
$\leq\|\tilde{f}\|_{L^{q_{0}}}^{\prod_{i=1}^{n}(1-1/r_{i})}$
$\leq\max\{\|\tilde{f}\|_{L^{q_{0}}},1\}.$
Hence,
$\int_{\mathbb{R}^{N}}|\tilde{f}(x)|^{q_{n}}\mathrm{d}x\leq\max\{\|\tilde{f}\|_{L^{q_{0}}},1\}^{q_{n}}.$
Letting $n\rightarrow\infty$, we see from Fatou’s lemma that $\tilde{f}\in L^{\tilde{q}}$. Consequently $f\in L^{\tilde{q}}$.
Observe that
$q_{n}=\frac{r_{n}N}{N-1}=\frac{N}{N-1}\Big{(}\frac{q_{n-1}}{{p^{\prime}_{1}}}+1\Big{)}.$
Letting $n\rightarrow\infty$, we get
$\tilde{q}=\frac{N}{N-1}\Big{(}\frac{\tilde{q}}{p^{\prime}_{1}}+1\Big{)}.$
Hence $1/p_{1}=1/\tilde{q}+1/N$. Therefore, $\tilde{q}=\tilde{p}_{0}$ and $f\in L^{\tilde{p}_{0}}$.
Set $r=\tilde{p}_{0}/p^{\prime}_{1}+1$. Then (2.10) turns out to be
$\|f\|_{L^{\tilde{p}_{0}}}\leq\|f\|_{L^{\tilde{p}_{0}}}^{1-1/r}\Big{(}Cr\|\nabla f\|_{L^{p_{1}}}\Big{)}^{1/r}.$
Finally, we have
$\|f\|_{L^{\tilde{p}_{0}}}\leq Cr\|\nabla f\|_{L^{p_{1}}}$
and this establishes (2.7) since $1/p_{1}-1/N=1/\tilde{q}=1/\tilde{p}_{0}$.
Moreover, (2.7) implies that $\|f\|_{L^{Np_{1}/(N-p_{1})}}\leq C\|f\|_{W_{1}^{\vec{p}}}$. Since $\|f\|_{L^{p_{0}}}\leq C^{\prime}\|f\|_{W_{1}^{\vec{p}}}$, by interpolation, we get that for any $q$ between $p_{0}$ and $Np_{1}/(N-p_{1})$,
$\|f\|_{L^{q}}\leq C\|f\|_{W_{1}^{\vec{p}}}.$
This completes the proof. ∎
The following is an immediate consequence.
Corollary 2.4.
Suppose that $\vec{p}=(p_{0},p_{1})$, $\vec{q}=(q_{0},q_{1})$, $1\leq p_{0},p_{1},q_{0},q_{1}<\infty$ and $p_{1}<N$. If $q_{1}=p_{1}$ and $q_{0}$ lies between $p_{0}$ and $Np_{1}/(N-p_{1})$, then
$W_{1}^{\vec{p}}(\mathbb{R}^{N})\hookrightarrow W_{1}^{\vec{q}}(\mathbb{R}^{N}).$
Remark 2.5.
The range for $q$ in the inequality (2.8) is the best possible, which can be shown as that for classical Sobolev spaces.
For example, take some $f\in W_{1}^{\vec{p}}$. Suppose that (2.8) is true for some $q$. Replacing $f(\cdot/\lambda)$ for $f$ in (2.8), we obtain
$\big{\|}f\big{(}\frac{\cdot}{\lambda}\big{)}\big{\|}_{L^{q}}\lesssim\big{\|}f\big{(}\frac{\cdot}{\lambda}\big{)}\big{\|}_{L^{p_{0}}}+\frac{1}{\lambda}\big{\|}\nabla f\big{(}\frac{\cdot}{\lambda}\big{)}\big{\|}_{L^{p_{1}}}.$
Thus we have
$\lambda^{N/q}\|f\|_{L^{q}}\lesssim\lambda^{N/p_{0}}\|f\|_{L^{p_{0}}}+\frac{1}{\lambda^{1-N/p_{1}}}\|\nabla f\|_{L^{p_{1}}}$
and therefore
$1\lesssim\lambda^{N(1/p_{0}-1/q)}+\lambda^{N(1/p_{1}-1/N-1/q)}.$
First, we assume that $p_{0}\leq Np_{1}/(N-p_{1})$. If $q>Np_{1}/(N-p_{1})$, then both $1/p_{1}-1/N-1/q$ and $1/p_{0}-1/q$ are positive. Letting $\lambda\rightarrow 0$, we get a contradiction.
If $q<p_{0}$, then both $1/p_{1}-1/N-1/q$ and $1/p_{0}-1/q$ are negative. Letting $\lambda\rightarrow\infty$, we also get a contradiction.
Next, we assume that $p_{0}>Np_{1}/(N-p_{1})$. With similar arguments we get a contradiction. Hence (2.8) is true if and only if $q$ is between $p_{0}$ and $Np_{1}/(N-p_{1})$.
With the help of Lemma 2.3, we prove the density of compactly supported infinitely many differentiable functions in nonuniform Sobolev spaces with general indices.
Theorem 2.6.
For any $k\geq 1$ and $\vec{p}\in[1,\infty)^{k+1}$, the space $C_{c}^{\infty}(\mathbb{R}^{N})$ is dense in $W_{k}^{\vec{p}}(\mathbb{R}^{N})$.
Proof.
By Lemma 2.2, it suffices to consider the case $1/p_{i_{0}}>1/p_{i_{0}-1}+1/N$ for some $i_{0}$ with $1\leq i_{0}\leq k$.
In this case, we have $N\geq 2$. Let $\vec{q}=(q_{0},\ldots,q_{N})$, where $q_{n}$ is defined recursively by
$q_{N}$ $=p_{N},$
$q_{n}$ =$=\begin{cases}p_{n},&\mbox{ if }1/p_{n}\geq 1/q_{n+1}-1/N,\\ Nq_{n+1}/(N-q_{n+1}),&\mbox{ otherwise},\end{cases}\qquad n=N-1,\ldots,0.$
By Lemma 2.3 (ii), we have $f\in W_{k}^{\vec{q}}$. Since $1/p_{n}\leq 1/q_{n}\leq 1/q_{n-1}+1/N$ for all $1\leq n\leq N$, we have
$\frac{1}{p_{n}}\leq\frac{1}{q_{n}}\leq\frac{1}{q_{m}}+\frac{n-m}{N},\quad n>m.$ (2.12)
It suffices to show that (2.5) is also true in this case.
In the sequel we adopt the notation introduced in the proof of Lemma 2.2. Recall that $\partial^{\gamma}f\in L^{q_{|\gamma|}}$. As in the proof of Lemma 2.2, we may assume that $\partial^{\gamma}g\in L^{r}$ for any $r>q_{|\gamma|}$. If $p_{|\alpha|}\geq q_{|\gamma|}$, then we get (2.5) as in the proof of Lemma 2.2. For the case $p_{|\alpha|}<q_{|\gamma|}$, we see from (2.12) that
$\frac{1}{p_{|\alpha|}}\leq\frac{1}{q_{|\gamma|}}+\frac{|\beta|}{N}.$
Hence there is some $r\geq N/|\beta|$ such that
$\frac{1}{p_{|\alpha|}}=\frac{1}{q_{|\gamma|}}+\frac{1}{r}.$
Applying Hölder’s inequality, we get
$\Big{\|}\frac{1}{n^{|\beta|}}\partial^{\beta}\psi(\frac{\cdot}{n})\partial^{\gamma}g\Big{\|}_{L^{p_{|\alpha|}}}\leq\frac{1}{n^{|\beta|-N/r}}\cdot\|\partial^{\beta}\psi\|_{L^{r}}\|\partial^{\gamma}g1_{\{|x|\geq n\}}\|_{L^{q_{|\gamma|}}}\rightarrow 0.$
This completes the proof. ∎
Below is an embedding theorem for higher-order nonuniform Sobolev spaces.
Theorem 2.7.
Let $k$ be a positive integer and $\vec{p}=(p_{0},\ldots,p_{k})\in[1,\infty)^{k+1}$.
  1. (i)
    If $kp_{k}<N$, then for any $q$ between $p_{0}$ and $Np_{k}/(N-kp_{k})$, we have
    $W_{k}^{\vec{p}}(\mathbb{R}^{N})\hookrightarrow L^{q}(\mathbb{R}^{N}).$
  2. (ii)
    If $kp_{k}=N$, then for any $q$ with $p_{0}\leq q<\infty$, we have
    $W_{k}^{\vec{p}}(\mathbb{R}^{N})\hookrightarrow L^{q}(\mathbb{R}^{N}).$
  3. (iii)
    If $kp_{k}>N$, then
    $W_{k}^{\vec{p}}(\mathbb{R}^{N})\hookrightarrow L^{\infty}(\mathbb{R}^{N}).$
Proof.
(i)   First, we consider the case $kp_{k}<N$. We see from Lemma 2.3 that the conclusion is true for $k=1$.
For the case $k\geq 2$, applying (2.7) recursively, we get
$\|f\|_{L^{Np_{k}/(N-kp_{k})}}$ $\leq C_{1}\sum_{|\alpha|=1}\|\partial^{\alpha}f\|_{L^{Np_{k}/(N-(k-1)p_{k})}}$
$\leq C_{2}\sum_{|\alpha|=2}\|\partial^{\alpha}f\|_{L^{Np_{k}/(N-(k-2)p_{k})}}$
[…]
$\leq C_{k}\sum_{|\alpha|=k}\|\partial^{\alpha}f\|_{L^{p_{k}}}.$ (2.13)
By interpolation, we get that for any $q$ between $p_{0}$ and $Np_{k}/(N-kp_{k})$, $W_{k}^{\vec{p}}\hookrightarrow L^{q}$.
(ii)   Next, we consider the case $kp_{k}=N$.
First, we assume that $k=1$ and $N>1$. Take some $f\in W_{1}^{\vec{p}}$, where $\vec{p}=(p_{0},N)$. Set $r=p_{0}(N-1)/N+1$ and $h=|f|^{r-1}f$. We have $\nabla h=r|f|^{r-1}\nabla f$. Since $(r-1)/p_{0}+1/N=1$, we see from Hölder’s inequality that
$\|\nabla h\|_{L^{1}}\leq r\|f\|_{L^{p_{0}}}^{r-1}\|\nabla f\|_{L^{N}}<\infty.$
Hence $\nabla h\in L^{1}$. By Lemma 2.3, $h\in L^{N/(N-1)}$. Therefore, $f\in L^{rN/(N-1)}=L^{p_{0}+N/(N-1)}$.
Replacing $p_{0}$ by $p_{0}+N/(N-1)$ in the preceding argument, we obtain that $f$ lies in $L^{p_{0}+2N/(N-1)}$. Repeating this procedure yields that $f\in L^{p_{0}+nN/(N-1)}$ for any $n\geq 1$. Hence $f\in L^{q}$ for any $q$ satisfying $p_{0}\leq q<\infty$.
For the case $k=N=1$, we see from (2.9) that for smooth functions $f$,
$\|f\|_{L^{\infty}}\lesssim\|f^{\prime}\|_{L^{1}}.$
Applying the density of $C_{c}(\mathbb{R})$ in $W_{1}^{\vec{p}}(\mathbb{R})$, we get that the above inequality is valid for all $f\in W_{1}^{\vec{p}}$. Hence $f\in L^{q}$ for any $q$ satisfying $p_{0}\leq q<\infty$.
Next we assume that $kp_{k}=N$ for some $k\geq 2$. Take some $f\in W_{k}^{\vec{p}}$, where $\vec{p}=(p_{0},\ldots,p_{k})$. For any multi-index $\alpha$ with $|\alpha|=1$, we have $\partial^{\alpha}f\in W_{k-1}^{(p_{1},\ldots,p_{k})}$. Note that $(k-1)p_{k}<N$. We see from (i) that $\partial^{\alpha}f\in L^{Np_{k}/(N-(k-1)p_{k})}=L^{N}$. Hence $f\in W_{1}^{(p_{0},N)}$. Now we see from arguments in the case $k=1$ that $f\in L^{q}$ for any $q$ satisfying $p_{0}\leq q<\infty$.
(iii)   Finally, we consider the case $kp_{k}>N$. We prove it by induction on $k$.
For $k=1$, the arguments in [16, pages 80-82] work well for the nonuniform case with minor changes. Specifically, applying the $L^{p_{1}}$ norm for derivatives and the $L^{p_{0}}$ norm for the function itself, the same arguments yield that $W_{1}^{\vec{p}}\hookrightarrow L^{\infty}$.
Now we assume that the conclusion is true for the cases $1$, $\ldots$, $k-1$. Consider the case $k$. Fix some $f\in W_{k}^{\vec{p}}$. There are two cases.
(a)   $(k-1)p_{k}\geq N$.
If $(k-1)p_{k}>N$, then we see from the inductive assumption that for any multi-index $\alpha$ with $|\alpha|=1$,
$\|\partial^{\alpha}f\|_{L^{\infty}}\lesssim\|\partial^{\alpha}f\|_{W_{k-1}^{(p_{1},\ldots,p_{k})}}\leq\|f\|_{W_{k}^{\vec{p}}}.$
If $(k-1)p_{k}=N$, then we see from (ii) that for any multi-index $\alpha$ with $|\alpha|=1$, $\partial^{\alpha}f\in L^{q}$ for all $q$ with $p_{0}\leq q<\infty$.
Hence for $(k-1)p_{k}\geq N$, there is some $q>\max\{N,p_{0}\}$ such that
$\|\nabla f\|_{L^{q}}\lesssim\|f\|_{W_{k}^{\vec{p}}}.$
Consequently, $f\in W_{1}^{(p_{0},q)}$. Applying the inductive assumption for the case $k=1$, we get
$\|f\|_{L^{\infty}}\lesssim\|f\|_{W_{1}^{(p_{0},q)}}\lesssim\|f\|_{W_{k}^{\vec{p}}}.$
(b)   $(k-1)p_{k}<N$.
In this case, we see from (i) that $f\in W_{1}^{(p_{0},q_{1})}$ with $q_{1}=Np_{k}/(N-(k-1)p_{k})$. Since $kp_{k}>N$, we have $q_{1}>N$. Applying the inductive assumption for the case $k=1$ again, we get
$\|f\|_{L^{\infty}}\lesssim\|f\|_{W_{1}^{(p_{0},q_{1})}}\lesssim\|f\|_{W_{k}^{\vec{p}}}.$
By induction, the embedding $\|f\|_{L^{\infty}}\lesssim\|f\|_{W_{k}^{\vec{p}}}$ is valid for all $k\geq 1$. ∎
Theorem 2.7 has some interesting consequences. For example, the following result shows that the Sobolev space $W_{k}^{\vec{q}}$ with $1\leq q_{k}<N/k$ and $1/q_{i}=1/q_{k}-(k-i)/N$ is the largest one among all Sobolev spaces $W_{k}^{\vec{p}}$ with $p_{k}=q_{k}$.
Corollary 2.8.
Suppose that $\vec{p}=(p_{0},\ldots,p_{k})\in[1,\infty)^{k+1}$ with $1\leq p_{k}<N/k$. Let $\vec{q}=(q_{0},\ldots,q_{k})$ be such that $q_{k}=p_{k}$ and $q_{i}=Np_{k}/(N-(k-i)p_{k})$ for $10\leq i\leq k-$ 1. Then we have
$W_{k}^{\vec{p}}\hookrightarrow W_{k}^{\vec{q}}.$
Moreover, set $\vec{q}^{(i)}=(q_{0},\ldots,q_{i})$, $1\leq i\leq k$. We have
$W_{k}^{\vec{q}^{(k)}}\hookrightarrow W_{k-1}^{\vec{q}^{(k-1)}}\hookrightarrow\cdots\hookrightarrow W_{1}^{\vec{q}^{(1)}}\hookrightarrow L^{q_{0}}.$
Proof.
We see from (2.13) that for any $f\in W_{k}^{\vec{p}}$,
$\|f\|_{W_{k}^{\vec{q}}}\leq C\sum_{|\alpha|=k}\|\partial^{\alpha}f\|_{L^{p_{k}}}\leq C\|f\|_{W_{k}^{\vec{p}}}.$
Hence $W_{k}^{\vec{p}}\hookrightarrow W_{k}^{\vec{q}}$. The second conclusion is obvious. This completes the proof. ∎
3. Nonuniform Fractional Sobolev Spaces
Recall that nonuniform fractional Sobolev spaces are introduced in Definition 1.2. We point out that for any $10<s<$ 1 and $\vec{p}=(p_{0},p_{1})\in[1,\infty)^{2}$, $W_{s}^{\vec{p}}(\mathbb{R}^{N})$ is a Banach space, a fact that can be proved with almost the same arguments as that used in the proof of [16, Proposition 4.24].
3.1. Embedding Theorem for Nonuniform Fractional Sobolev Spaces
The main result in this subsection is the following embedding theorem.
Theorem 3.1.
Suppose that $\vec{p}=(p_{0},p_{1})$ with $1\leq p_{0}<\infty$, $1<p_{1}<\infty$ and $10<s<$ 1. When $p_{0}\leq p_{1}$, we have
  1. (i)
    If $sp_{1}<N$, then $W_{s}^{\vec{p}}(\mathbb{R}^{N})\hookrightarrow L^{q}(\mathbb{R}^{N})$ for all $p_{0}\leq q\leq Np_{1}/(N-sp_{1})$.
  2. (ii)
    If $sp_{1}=N$, then $W_{s}^{\vec{p}}(\mathbb{R}^{N})\hookrightarrow L^{q}(\mathbb{R}^{N})$ for all $p_{0}\leq q<\infty$.
  3. (iii)
    If $sp_{1}>N$, then $W_{s}^{\vec{p}}(\mathbb{R}^{N})\hookrightarrow L^{\infty}(\mathbb{R}^{N})$.
For the case $p_{0}>p_{1}$, if $sp_{1}<N$ and $p_{0}<Np_{1}/(N-sp_{1})$, we also have $W_{s}^{\vec{p}}(\mathbb{R}^{N})\hookrightarrow L^{q}(\mathbb{R}^{N})$ for all $p_{0}\leq q\leq Np_{1}/(N-sp_{1})$.
To prove Theorem 3.1, we need some preliminary results.
We see from [16, Lemma 4.33] that for any $f\in L^{p}$,
$\iint_{\mathbb{R}^{N}\times\mathbb{R}^{N}}\frac{|f(x)-f(y)|^{p}}{|x-y|^{N+ps}}\mathrm{d}x\,\mathrm{d}y<\infty$
is equivalent to
[…]
Checking the arguments in the proof of [16, Lemma 4.33], we find that the hypothesis $f\in L^{p}$ is not used. Moreover, with the same arguments we get the following result.
Proposition 3.2.
Let $10<s<$ 1 and $1\leq p<\infty$. Then there exist positive constants $C_{1}$ and $C_{2}$ such that for any measurable function $f$ which is finite almost everywhere,
[…]
The following construction is used in the proof of the embedding theorem. Unlike the classical case, here we need that $\varphi$ equals $1$ in a neighborhood of 00 to ensure that its derivatives vanish near 00.
Lemma 3.3.
Suppose that $1\leq p_{0},p_{1}<\infty$, $0<s\leq 1$, $f\in W_{s}^{\vec{p}}(\mathbb{R}^{N})$, $\varphi\in C_{c}^{\infty}(\mathbb{R})$ with $\varphi(t)=1$ for $|t|<1$ and $\varphi(t)=0$ for $|t|>A$, where $A$ is a constant. Let $\alpha=1-1/p_{1}-s$ for $10<s<$ 1 and $\alpha=-1/p_{1}+\eta$ for $s=1$, where $10<\eta<$ 1. Set
$g(t,x)$ $=\frac{\varphi(t)}{t^{N}}\int_{[0,t]^{N}}f(x+y)\,\mathrm{d}y,\quad t>0,$ (3.1)
$g(0,x)$ $=\lim_{t\rightarrow 0}g(t,x).$ (3.2)
We have
$\left\|t^{\alpha}\nabla_{x}g(t,x)\right\|_{L^{p_{1}}(\mathbb{R}^{N+1})}$ $\lesssim\|f\|_{W_{s}^{\vec{p}}}.$ (3.3)
Moreover, there are two functions $h_{0}$ and $h_{1}$ such that $\partial_{t}g(t,x)=h_{0}(t,x)+h_{1}(t,x)$ and
$\left\|t^{\alpha}h_{0}(t,x)\right\|_{L^{p_{1}}(\mathbb{R}^{N+1})}$ $\lesssim\|f\|_{W_{s}^{\vec{p}}},$ (3.4)
$\left\|h_{1}(t,\cdot)\right\|_{L^{p_{0}}(\mathbb{R}^{N})}$ $\lesssim\|f\|_{L^{p_{0}}}1_{[1,A]}(t).$ (3.5)
Furthermore, if $p_{0}\leq p_{1}$, we have
$\left\|t^{\alpha}\nabla g(t,x)\right\|_{L^{p_{1}}(\mathbb{R}^{N+1})}$ $\lesssim\|f\|_{W_{s}^{\vec{p}}}.$ (3.6)
Proof.
First, we estimate $\|\partial_{x_{i}}g(t,x)\|_{L^{p_{1}}(\mathbb{R}^{N+1})}$, $1\leq i\leq N$. Denote
$\hat{x}_{i}$ $=x-x_{i}e_{i}\mbox{\quad and\quad}\mathrm{d}\hat{x}_{i}=\prod_{j\neq i}\mathrm{d}x_{j}.$
Observe
$g(t,x)$ $=\frac{\varphi(t)}{t^{N}}\int_{x_{i}}^{x_{i}+t}\int_{[0,t]^{N-1}}f(\hat{x}_{i}+\hat{y}_{i}+y_{i}e_{i})\,\mathrm{d}\hat{y}_{i}\,\mathrm{d}y_{i}.$
We get
$\partial_{x_{i}}g(t,x)$ $=\frac{\varphi(t)}{t^{N}}\int_{[0,t]^{N-1}}\Big{(}f(\hat{x}_{i}+\hat{y}_{i}+(x_{i}+t)e_{i})-f(\hat{x}_{i}+\hat{y}_{i}+x_{i}e_{i})\Big{)}\mathrm{d}\hat{y}_{i}$
$=\frac{\varphi(t)}{t^{N}}\int_{[0,t]^{N-1}}\Big{(}f(x+te_{i}+\hat{y}_{i})-f(x+\hat{y}_{i})\Big{)}\mathrm{d}\hat{y}_{i}$
$=\frac{\varphi(t)}{t}\int_{[0,1]^{N-1}}\Big{(}f(x+te_{i}+t\hat{y}_{i})-f(x+t\hat{y}_{i})\Big{)}\mathrm{d}\hat{y}_{i}.$
Hence for $10<s<$ 1,
$\|t^{\alpha}\partial_{x_{i}}g(t,x)\|_{L^{p_{1}}(\mathbb{R}^{N+1})}^{p_{1}}\!$ $\leq\!\|\varphi(t)\|_{L^{\infty}}^{p_{1}}\int_{0}^{\infty}\!\!\!\int_{\mathbb{R}^{N}}\!\!\int_{[0,1]^{N-1}}\!\!\!\frac{|f(x+te_{i}+t\hat{y}_{i})-f(x+t\hat{y}_{i})|^{p_{1}}}{t^{1+sp_{1}}}\mathrm{d}\hat{y}_{i}\,\mathrm{d}x\,\mathrm{d}t$
$\lesssim[f]_{W_{s}^{p_{1}}}^{p_{1}}.$
For the case $s=1$, we have $\alpha=-1/p_{1}+\eta$. We see from (3.1) that
$\nabla_{x}g(t,x)=\frac{\varphi(t)}{t^{N}}\int_{[0,t]^{N}}\nabla f(x+y)\mathrm{d}y.$
Thus we obtain
$\|t^{\alpha}\nabla_{x}g(t,x)\|_{L^{p_{1}}(\mathbb{R}^{N+1})}\leq\|t^{\alpha}\varphi(t)\|_{L^{p_{1}}}\|\nabla f\|_{L^{p_{1}}}<\infty$
proving (3.3).
Next we deal with $\partial_{t}g(t,x)$. A simple computation shows that
$\partial_{t}g(t,x)$ $=\varphi(t)\bigg{(}\frac{-N}{t^{N+1}}\int_{[0,t]^{N}}f(x+y)\mathrm{d}y+\frac{1}{t^{N}}\sum_{i=1}^{N}\int_{[0,t]^{N-1}}f(x+\hat{y}_{i}+te_{i})\mathrm{d}\hat{y}_{i}\bigg{)}$
$\qquad+\frac{\varphi^{\prime}(t)}{t^{N}}\int_{[0,t]^{N}}f(x+y)\mathrm{d}y$
$=\frac{\varphi(t)}{t^{N+1}}\sum_{i=1}^{N}\int_{[0,t]^{N}}\Big{(}f(x+\hat{y}_{i}+te_{i})-f(x+y)\Big{)}\mathrm{d}y+\varphi^{\prime}(t)\int_{[0,1]^{N}}f(x+ty)\mathrm{d}y$
$=\frac{\varphi(t)}{t}\sum_{i=1}^{N}\int_{[0,1]^{N}}\Big{(}f(x+t\hat{y}_{i}+te_{i})-f(x+ty)\Big{)}\mathrm{d}y+\varphi^{\prime}(t)\int_{[0,1]^{N}}f(x+ty)\mathrm{d}y$
$:=h_{0}(t,x)+h_{1}(t,x).$
Next, we prove (3.4). For $10<s<$ 1, we have
$\|t^{\alpha}h_{0}(t,x)\|_{L^{p_{1}}(\mathbb{R}^{N+1})}^{p_{1}}$ $\lesssim\|\varphi\|_{L^{\infty}}^{p_{1}}\sum_{i=1}^{N}\int_{0}^{\infty}\!\int_{[0,1]^{N}}\!\int_{\mathbb{R}^{N}}\frac{|f(x+t\hat{y}_{i}+te_{i})-f(x+ty)|^{p_{1}}}{t^{1+sp_{1}}}\mathrm{d}x\,\mathrm{d}y\,\mathrm{d}t.$
By the change of variables $x\rightarrow x+ty$, we write
$\|t^{\alpha}h_{0}(t,x)\|_{L^{p_{1}}(\mathbb{R}^{N+1})}^{p_{1}}$ $\lesssim\|\varphi\|_{L^{\infty}}^{p_{1}}\sum_{i=1}^{N}\int_{0}^{\infty}\int_{[0,1]^{N}}\int_{\mathbb{R}^{N}}\frac{|f(x+t(1-y_{i})e_{i})-f(x)|^{p_{1}}}{t^{1+sp_{1}}}\mathrm{d}x\,\mathrm{d}y\,\mathrm{d}t$
$=\|\varphi\|_{L^{\infty}}^{p_{1}}\sum_{i=1}^{N}\int_{0}^{\infty}\int_{0}^{1}(1-y_{i})^{sp_{1}}\int_{\mathbb{R}^{N}}\frac{|f(x+te_{i})-f(x)|^{p_{1}}}{t^{1+sp_{1}}}\mathrm{d}x\,\mathrm{d}y_{i}\,\mathrm{d}t$
$\lesssim[f]_{W_{s}^{p_{1}}}^{p_{1}},$ (3.7)
having used the change of variables $t\to(1-y_{i})t$. And for the case $s=1$, we have
$\|t^{\alpha}h_{0}(t,x)\|_{L_{x}^{p_{1}}}$ $\leq|t^{\alpha-1}\varphi(t)|\sum_{i=1}^{N}\int_{[0,1]^{N}}\Big{\|}f(x+t\hat{y}_{i}+te_{i})-f(x+ty)\Big{\|}_{L_{x}^{p_{1}}}\mathrm{d}y$
$=|t^{\alpha-1}\varphi(t)|\sum_{i=1}^{N}\int_{[0,1]^{N}}\Big{\|}\int_{ty_{i}}^{t}\partial_{x_{i}}f(x+t\hat{y}_{i}+\tau e_{i})\mathrm{d}\tau\Big{\|}_{L_{x}^{p_{1}}}\mathrm{d}y$
$\lesssim|t^{\alpha}\varphi(t)|\cdot\|\nabla f\|_{L^{p_{1}}}.$
Hence
$\|t^{\alpha}h_{0}(t,x)\|_{L^{p_{1}}(\mathbb{R}^{N+1})}\lesssim\|\nabla f\|_{L^{p_{1}}}.$ (3.8)
To prove (3.5), note that $\varphi^{\prime}(t)=0$ for $10<t<$ 1 or $t>A$. For $1\leq t\leq A$, we see from Minkowski’s inequality that
$\|h_{1}(t,\cdot)\|_{L^{p_{0}}(\mathbb{R}^{N})}$ $\leq|\varphi^{\prime}(t)|\int_{[0,1]^{N}}\|f(\cdot+ty)\|_{L^{p_{0}}}\mathrm{d}y\leq|\varphi^{\prime}(t)|\cdot\|f\|_{L^{p_{0}}}.$ (3.9)
Hence (3.5) is valid.
Finally, we consider the case $p_{0}\leq p_{1}$. To prove (3.6), it suffices to show that $\|t^{\alpha}h_{1}(t,x)\|_{L^{p_{1}}(\mathbb{R}^{N+1})}\lesssim\|f\|_{L^{p_{0}}}$, thanks to (3.3) and (3.4).
Recall that $\varphi^{\prime}(t)=0$ for $10<t<$ 1. For $t\geq 1$, we have
$|h_{1}(t,x)|$ $\leq|\varphi^{\prime}(t)|\Big{(}\int_{[0,1]^{N}}|f(x+ty)|^{p_{0}}\mathrm{d}y\Big{)}^{1/p_{0}}$
$=|\varphi^{\prime}(t)|\Big{(}\frac{1}{t^{N}}\int_{[0,t]^{N}}|f(x+y)|^{p_{0}}\mathrm{d}y\Big{)}^{1/p_{0}}$
$\leq|\varphi^{\prime}(t)|\cdot\|f\|_{L^{p_{0}}}.$
Hence
$\|t^{\alpha}h_{1}(t,x)\|_{L^{p_{1}}(\mathbb{R}^{N+1})}^{p_{1}}$ $\leq\int_{\mathbb{R}^{N}}\int_{1}^{\infty}|t^{\alpha}\varphi^{\prime}(t)|^{p_{1}}\|f\|_{L^{p_{0}}}^{p_{1}-p_{0}}\Big{(}\int_{[0,1]^{N}}|f(x+ty)|\mathrm{d}y\Big{)}^{p_{0}}\,\mathrm{d}t\,\mathrm{d}x$
$\leq\int_{\mathbb{R}^{N}}\int_{1}^{\infty}|t^{\alpha}\varphi^{\prime}(t)|^{p_{1}}\|f\|_{L^{p_{0}}}^{p_{1}-p_{0}}\int_{[0,1]^{N}}|f(x+ty)|^{p_{0}}\,\mathrm{d}y\,\mathrm{d}t\,\mathrm{d}x$
$\leq\|f\|_{L^{p_{0}}}^{p_{1}}.$
This completes the proof. ∎
For the case $1\leq p_{0}\leq p_{1}<\infty$, the nonuniform Sobolev space $W_{s}^{(p_{0},p_{1})}$ is a subspace of the classical one $W_{s}^{p_{1}}$. Moreover, the inclusion is in fact an embedding and is also valid for $s=1$.
Lemma 3.4.
Suppose that $0<s\leq 1$ and $1\leq p_{0}\leq p_{1}<\infty$. We have
$W_{s}^{\vec{p}}\hookrightarrow W_{s}^{p_{1}}.$
Moreover, the inclusion is not true if $p_{0}>p_{1}$.
Proof.
Let $\varphi$ and $\psi$ be functions in $C_{c}^{\infty}(\mathbb{R}^{N})$ and $C_{c}^{\infty}(\mathbb{R})$, respectively, with values between 00 and $1$ and equal to $1$ in neighborhoods of 00. Set $\Phi(t,x)=\varphi(x)\psi(t)$. Take some $A>0$ such that $\mathrm{supp}\,\Phi\subset[-A,A]^{N+1}$.
Fix some $f\in W_{s}^{\vec{p}}$. Let $\alpha$ and $g$ be defined as in Lemma 3.3, $\delta_{0}$ be the Dirac measure at 00, and
$E(t,x)=\begin{cases}\dfrac{c_{N}}{(t^{2}+|x|^{2})^{(N-1)/2}},&N\geq 2,\\ c_{1}\log(t^{2}+|x|^{2}),&N=1,\end{cases}$
be the fundamental solution of the Laplacian satisfying $\Delta E=\delta_{0}$.
Recall that for a tempered distribution $\Lambda$ and a function $h\in C^{\infty}$, for which the function itself and all of its derivatives have at most polynomial growth at infinity, the product $h\Lambda$ is the tempered distribution [21, Definition 2.3.15] defined by
$\langle h\Lambda,u\rangle=\langle\Lambda,hu\rangle,\qquad\forall u\in\mathscr{S}.$ (3.10)
Since $\Phi(t,x)$ equals $1$ in a neighbourhood of 00, we have
$\Phi\Delta E=\Phi\delta_{0}=\delta_{0}.$
Hence
$g$ $=\delta_{0}*g$
$=\Delta(\Phi E)*g-2(\nabla\Phi\cdot\nabla E)*g-((\Delta\Phi)E)*g$
$=((\nabla\Phi)E)*\nabla g+(\Phi\nabla E)*\nabla g-2(\nabla\Phi\cdot\nabla E)*g-((\Delta\Phi)E)*g,$ (3.11)
where we use the notation
$u*v=\sum_{i=1}^{N+1}u_{i}*v_{i}$
to denote the convolution of two $\mathbb{C}^{N+1}$-valued functions.
Since $g(0,x)=f(x)$ a.e., it suffices to show that $\|g(0,\cdot)\|_{L^{p_{1}}}\lesssim\|f\|_{W_{s}^{\vec{p}}}$. We prove it in the following steps.
(S1)  First, we estimate $\|((\nabla\Phi)E)*\nabla g(0,\cdot)\|_{L^{p_{1}}}$. For $N\geq 2$, we have
$((\nabla\Phi)E)*\nabla g(0,x)=\int_{0}^{A}\int_{\mathbb{R}^{N}}\frac{(\nabla\Phi)(-t,y)\cdot\nabla g(t,x-y)}{(t^{2}+|y|^{2})^{(N-1)/2}}\mathrm{d}y\,\mathrm{d}t.$
Applying Minkovski’s inequality, we get
$\|((\nabla\Phi)E)*\nabla g(0,\cdot)\|_{L^{p_{1}}}$ $\leq\|\nabla\Phi\|_{L^{\infty}}\int_{0}^{A}\|\nabla g(t,\cdot)\|_{L^{p_{1}}}\int_{[-A,A]^{N}}\frac{\mathrm{d}y}{(t+|y|)^{N-1}}\,\mathrm{d}t$
$\lesssim\int_{0}^{A}\|\nabla g(t,\cdot)\|_{L^{p_{1}}}\int_{[0,A]^{2}}\frac{\mathrm{d}y_{1}\mathrm{d}y_{2}}{t+y_{1}+y_{2}}\,\mathrm{d}t$
$\approx\int_{0}^{A}\|\nabla g(t,\cdot)\|_{L^{p_{1}}}\int_{0}^{A}\log\frac{t+y_{1}+A}{t+y_{1}}\mathrm{d}y_{1}\,\mathrm{d}t$
$\lesssim\int_{0}^{A}\|\nabla g(t,\cdot)\|_{L^{p_{1}}}\int_{0}^{A}\frac{1}{(t+y_{1})^{\varepsilon}}\mathrm{d}y_{1}\,\mathrm{d}t\mbox{\qquad$(0<\varepsilon<1)$}$
$\lesssim\int_{0}^{A}\|\nabla g(t,\cdot)\|_{L^{p_{1}}}\mathrm{d}t$
$\lesssim\|t^{-\alpha}\cdot 1_{[0,A]}(t)\|_{L^{p^{\prime}_{1}}}\|t^{\alpha}\nabla g(t,x)\|_{L^{p_{1}}(\mathbb{R}^{N+1})}$
$\approx\|t^{\alpha}\nabla g(t,x)\|_{L^{p_{1}}(\mathbb{R}^{N+1})}$
$\lesssim\|f\|_{W_{s}^{\vec{p}}},$
where we applied (3.6) in the last step. When $N\geq 3$, the first inequality is obtained by successively integrating the variables $y_{N},y_{N-1},\dots,y_{3}$ over $[0,\infty)$; the effect of this is the reduction of the exponent by $1$ in each integration, so after $N-2$ integrations, the exponent becomes $(N-1)-(N-2)=1$.
For the case $N=1$, we have $|E(t,y)|=|c_{N}\log(t^{2}+y^{2})|\lesssim(|t|+|y|)^{-\varepsilon}$ for any $\varepsilon>0$. Hence
$\|((\nabla\Phi)E)*\nabla g(0,\cdot)\|_{L^{p_{1}}}$ $\lesssim\int_{0}^{A}\|\nabla g(t,\cdot)\|_{L^{p_{1}}}\int_{0}^{A}\frac{1}{(t+y_{1})^{\varepsilon}}\mathrm{d}y_{1}\,\mathrm{d}t\mbox{\qquad$(0<\varepsilon<1)$}$
$\lesssim\|f\|_{W_{s}^{\vec{p}}}.$
(S2)  Next, we estimate $\|(\Phi\nabla E)*\nabla g(0,\cdot)\|_{L^{p_{1}}}$. Observe that for $(t,y)\in[-A,A]^{N+1}$,
$|\nabla E(t,y)|\lesssim\frac{1}{(|t|+|y|)^{N}}.$
Similar arguments as the previous case show that
$\|(\Phi\nabla E)*\nabla g(0,\cdot)\|_{L^{p_{1}}}$ $\lesssim\int_{0}^{A}\|\nabla g(t,\cdot)\|_{L^{p_{1}}}\int_{[-A,A]^{N}}\frac{\mathrm{d}y}{(t+|y|)^{N}}\,\mathrm{d}t$
$\approx\int_{0}^{A}\|\nabla g(t,\cdot)\|_{L^{p_{1}}}\int_{0}^{A}\frac{\mathrm{d}y_{1}}{t+y_{1}}\,\mathrm{d}t$
$\approx\int_{0}^{A}\|\nabla g(t,\cdot)\|_{L^{p_{1}}}\log\frac{t+A}{t}\mathrm{d}t$
$\lesssim\int_{0}^{A}\|\nabla g(t,\cdot)\|_{L^{p_{1}}}t^{-\varepsilon}\mathrm{d}t$
$\lesssim\|t^{-(\alpha+\varepsilon)}\cdot 1_{[0,A]}(t)\|_{L^{p^{\prime}_{1}}}\|t^{\alpha}\nabla g(t,x)\|_{L^{p_{1}}(\mathbb{R}^{N+1})}$
$\lesssim\|f\|_{W_{s}^{\vec{p}}},$
where $0<\varepsilon<s$ for $s<1$ and $0<\varepsilon<1-\eta$ for $s=1$. Again, we apply (3.6) in the last step.
(S3)  Now we deal with the last two terms.
Observe that both $\nabla\Phi$ and $\Delta\Phi$ vanish in a neighborhood of 00. Hence both $(\nabla\Phi\cdot\nabla E)$ and $(\Delta\Phi)E$ are compactly supported bounded functions. Hence
$F(x):=|2(\nabla\Phi\cdot\nabla E)*g(0,x)|+|((\Delta\Phi)E)*g(0,x)|$ $\lesssim\int_{0}^{A}\int_{[-A,A]^{N}}|g(t,x-y)|\mathrm{d}y\,\mathrm{d}t.$
For any $q>p_{0}$, there is some $r>1$ such that
$\frac{1}{r}+\frac{1}{p_{0}}=\frac{1}{q}+1.$
Observe that
$\|g(t,\cdot)\|_{L^{p_{0}}}=|\varphi(t)|\cdot\Big{\|}\int_{[0,1]^{N}}f(\cdot+ty)\mathrm{d}y\Big{\|}_{L^{p_{0}}}\lesssim\|\varphi\|_{L^{\infty}}\cdot\|f\|_{L^{p_{0}}}.$ (3.12)
We see from Young’s inequality that
$\|F\|_{L^{q}}$ $\lesssim\|f\|_{L^{p_{0}}}.$
Setting $q=p_{1}$, we get $\|F\|_{L^{p_{1}}}\lesssim\|f\|_{L^{p_{0}}}$.
Combing results in (S1), (S2), and (S3), yields
$W_{s}^{(p_{0},p_{1})}\hookrightarrow W_{s}^{(p_{1},p_{1})}=W_{s}^{p_{1}}.$
For the case $p_{0}>p_{1}$, we provide a counterexample in Example 3.6. This completes the proof. ∎
The following is an immediate consequence.
Corollary 3.5.
Suppose that $\vec{p}=(p_{0},p_{1})$ with $1\leq p_{0}\leq p_{1}<\infty$. For any $10<\tilde{s}<s<$ 1, we have
$W_{1}^{\vec{p}}\hookrightarrow W_{s}^{\vec{p}}\hookrightarrow W_{\tilde{s}}^{\vec{p}}.$ (3.13)
Proof.
First, we prove that $W_{1}^{\vec{p}}\hookrightarrow W_{s}^{\vec{p}}$. Take some $f\in W_{1}^{\vec{p}}$. As for functions in classical Sobolev spaces, for almost all $x\in\mathbb{R}^{N}$, we have
$|f(x)-f(x+ae_{j})|^{p_{1}}=\Big{|}\int_{0}^{a}\partial_{x_{j}}f(x+te_{j})\mathrm{d}t\Big{|}^{p_{1}}\leq|a|^{p_{1}-1}\int_{0}^{a}|\partial_{x_{j}}f(x+te_{j})|^{p_{1}}\mathrm{d}t.$
Hence
$\int_{\mathbb{R}^{N}}|f(x)-f(x+ae_{j})|^{p_{1}}\mathrm{d}x\leq|a|^{p_{1}}\|\partial_{x_{j}}f\|_{L^{p_{1}}}^{p_{1}}.$
Therefore,
$\int_{|a|<1}\int_{\mathbb{R}^{N}}\frac{|f(x)-f(x+ae_{j})|^{p_{1}}}{|a|^{1+sp_{1}}}\mathrm{d}x\,\mathrm{d}a\leq\int_{|a|<1}\frac{|a|^{p_{1}}}{|a|^{1+sp_{1}}}\|\partial_{x_{j}}f\|_{L^{p_{1}}}^{p_{1}}\mathrm{d}a\lesssim\|\partial_{x_{j}}f\|_{L^{p_{1}}}^{p_{1}}.$
On the other hand, since $p_{0}\leq p_{1}$, we have $W_{1}^{\vec{p}}\hookrightarrow W_{1}^{p_{1}}$, thanks to Lemma 3.4. Hence
$\int_{|a|\geq 1}\int_{\mathbb{R}^{N}}\frac{|f(x)-f(x+ae_{j})|^{p_{1}}}{|a|^{1+sp_{1}}}\mathrm{d}x\,\mathrm{d}a\lesssim\int_{|a|\geq 1}\frac{\|f\|_{L^{p_{1}}}^{p_{1}}}{|a|^{1+sp_{1}}}\mathrm{d}a\lesssim\|f\|_{L^{p_{1}}}^{p_{1}}.$
Consequently, $\|f\|_{W_{s}^{\vec{p}}}\lesssim\|f\|_{W_{1}^{\vec{p}}}$.
Next, we show that $W_{s}^{\vec{p}}\hookrightarrow W_{\tilde{s}}^{\vec{p}}$. Fix some $f\in W_{s}^{\vec{p}}$. Since $\tilde{s}<s$, we have
$\int_{|a|<1}\int_{\mathbb{R}^{N}}\frac{|f(x)-f(x+ae_{j})|^{p_{1}}}{a^{1+p_{1}\tilde{s}}}\mathrm{d}x\,\mathrm{d}a\leq\int_{|a|<1}\int_{\mathbb{R}^{N}}\frac{|f(x)-f(x+ae_{j})|^{p_{1}}}{a^{1+p_{1}s}}\mathrm{d}x\,\mathrm{d}a\lesssim[f]_{W_{s}^{p_{1}}}^{p_{1}}.$
On the other hand, by Lemma 3.4, $f\in L^{p_{1}}$. Hence
$\int_{|a|\geq 1}\int_{\mathbb{R}^{N}}\frac{|f(x)-f(x+ae_{j})|^{p_{1}}}{a^{1+p_{1}\tilde{s}}}\mathrm{d}x\,\mathrm{d}a\leq\int_{|a|\geq 1}\frac{C_{p_{1}}\|f\|_{L^{p_{1}}}^{p_{1}}}{a^{1+p_{1}\tilde{s}}}\mathrm{d}a\lesssim\|f\|_{L^{p_{1}}}^{p_{1}}.$
This completes the proof. ∎
We point out that the inclusion in Corollary 3.5 is not true in general for the case $p_{0}>p_{1}$, which is quite different from the classical case, since for any $1\leq p<\infty$ and $10<\tilde{s}<s<$ 1,
$W_{1}^{p}\hookrightarrow W_{s}^{p}\hookrightarrow W_{\tilde{s}}^{p}.$
Below is a counterexample.
Example 3.6.
Suppose that $0<s\leq 1$, $1\leq p_{0},p_{1}<\infty$ and $p_{1}\alpha<N<p_{0}\alpha$. Let $f(x)=(1+|x|^{2})^{-\alpha/2}$, $x\in\mathbb{R}^{N}$ and let $\vec{p}=(p_{0},p_{1})$. We have $f\in W_{s}^{\vec{p}}(\mathbb{R}^{N})$ if and only if $p_{1}(\alpha+s)>N$.
Consequently, for any $0<s\leq 1$ and $p_{0}>p_{1}$, $W_{s}^{\vec{p}}\not\subset W_{s}^{p_{1}}$.
Moreover, if $0<\tilde{s}<s\leq 1$ and $\tilde{s}<N/p_{1}-N/p_{0}$, we have $W_{s}^{\vec{p}}\not\subset W_{\tilde{s}}^{\vec{p}}$.
Proof.
(i)  We show that $f\in W_{s}^{\vec{p}}(\mathbb{R}^{N})$ if and only if $p_{1}(\alpha+s)>N$, for which we only need to consider the case $10<s<$ 1 since the other case $s=1$ is obvious.
First, we assume that $p_{1}(\alpha+s)>N$. Let us estimate the integral
$\int_{\mathbb{R}^{N}}|f(x)-f(x+ae_{j})|^{p_{1}}\mathrm{d}x,\qquad 1\leq j\leq N.$
Denote $\hat{x}_{j}=x-x_{j}e_{j}$ and $\mathrm{d}\hat{x}_{j}=\prod_{i\neq j}\mathrm{d}x_{i}$. Observe that
$|f(x)-f(x+ae_{j})|^{p_{1}}=\Big{|}\int_{0}^{a}\partial_{x_{j}}f(x+te_{j})\mathrm{d}t\Big{|}^{p_{1}}\leq|a|^{p_{1}-1}\Big{|}\int_{0}^{a}|\partial_{x_{j}}f(x+te_{j})|^{p_{1}}\mathrm{d}t\Big{|}.$ (3.14)
When $|a|>1$, $|t|\leq|a|$ and $|x|>2|a|$, we have $|x+te_{j}|>|a|$. Hence
$\int_{|x|>2|a|}|f(x)-f(x+ae_{j})|^{p_{1}}\mathrm{d}x$
$\lesssim|a|^{p_{1}-1}\bigg{|}\int_{0}^{a}\int_{|x|>2|a|}\frac{|x_{j}+t|^{p_{1}}}{(1+|\hat{x}_{j}|^{2}+|x_{j}+t|^{2})^{(\alpha+2)p_{1}/2}}\mathrm{d}x\,\mathrm{d}t\bigg{|}$
$\leq|a|^{p_{1}-1}\bigg{|}\int_{0}^{a}\int_{|x|>|a|}\frac{|x_{j}|^{p_{1}}}{(1+|\hat{x}_{j}|^{2}+|x_{j}|^{2})^{(\alpha+2)p_{1}/2}}\mathrm{d}x\,\mathrm{d}t\bigg{|}$
$=|a|^{p_{1}}\bigg{(}\int_{|x_{j}|>|a|}\int_{\mathbb{R}^{N-1}}\frac{|x_{j}|^{p_{1}}}{(1+|\hat{x}_{j}|^{2}+|x_{j}|^{2})^{(\alpha+2)p_{1}/2}}\mathrm{d}\hat{x}_{j}\,\mathrm{d}x_{j}$
$\qquad+\int_{|x_{j}|<|a|}\int_{|\hat{x}_{j}|>(a^{2}-x_{j}^{2})^{1/2}}\frac{|x_{j}|^{p_{1}}}{(1+|\hat{x}_{j}|^{2}+|x_{j}|^{2})^{(\alpha+2)p_{1}/2}}\mathrm{d}\hat{x}_{j}\,\mathrm{d}x_{j}\bigg{)}$
$\approx|a|^{p_{1}}\bigg{(}\int_{|x_{j}|>|a|}\frac{|x_{j}|^{p_{1}}\mathrm{d}x_{j}}{(1+|x_{j}|)^{(\alpha+2)p_{1}-(N-1)}}+\int_{|x_{j}|<|a|}\frac{|x_{j}|^{p_{1}}\mathrm{d}x_{j}}{(1+|a|)^{(\alpha+2)p_{1}-(N-1)}}\bigg{)}$
$\approx|a|^{N-p_{1}\alpha},$ (3.15)
where we use the fact that $(\alpha+1)p_{1}>(\alpha+s)p_{1}>N$. On the other hand,
$\int_{|x|\leq 2|a|}|f(x)-f(x+ae_{j})|^{p_{1}}\mathrm{d}x$
$\lesssim\int_{|x|\leq 2|a|}\Big{(}\frac{1}{(1+|x|^{2})^{p_{1}\alpha/2}}+\frac{1}{(1+|x+ae_{j}|^{2})^{p_{1}\alpha/2}}\Big{)}\mathrm{d}x$
$\leq\int_{|x|\leq 2|a|}\frac{1}{(1+|x|^{2})^{p_{1}\alpha/2}}\mathrm{d}x+\int_{|x+ae_{j}|\leq 3|a|}\frac{1}{(1+|x+ae_{j}|^{2})^{p_{1}\alpha/2}}\mathrm{d}x$
$\lesssim|a|^{N-p_{1}\alpha}.$ (3.16)
Putting (3.15) and (3.16) together, we get
$\int_{\mathbb{R}^{N}}|f(x)-f(x+ae_{j})|^{p_{1}}\mathrm{d}x\lesssim|a|^{N-p_{1}\alpha},\quad|a|>1.$
For $|a|\leq 1$, since $|\partial_{x_{j}}f(x)|\lesssim 1/(1+|x|^{2})^{(\alpha+1)/2}$ and $p_{1}(\alpha+1)>p_{1}(\alpha+s)>N$, we have $\partial_{x_{j}}f\in L^{p_{1}}$. It follows from (3.14) that
$\int_{\mathbb{R}^{N}}|f(x)-f(x+ae_{j})|^{p_{1}}\mathrm{d}x\leq|a|^{p_{1}}\|\partial_{x_{j}}f\|_{L^{p_{1}}}^{p_{1}}.$
Combining the preceding estimates we deduce
$\int_{\mathbb{R}}\int_{\mathbb{R}^{N}}\frac{|f(x)-f(x+ae_{j})|^{p_{1}}}{|a|^{1+p_{1}s}}\mathrm{d}x\,\mathrm{d}a$ $=\int_{|a|\leq 1}\int_{\mathbb{R}^{N}}\frac{|f(x)-f(x+ae_{j})|^{p_{1}}}{|a|^{1+p_{1}s}}\mathrm{d}x\,\mathrm{d}a$
$\qquad+\int_{|a|>1}\int_{\mathbb{R}^{N}}\frac{|f(x)-f(x+ae_{j})|^{p_{1}}}{|a|^{1+p_{1}s}}\mathrm{d}x\,\mathrm{d}a$
$\lesssim\int_{|a|\leq 1}|a|^{p_{1}-1-p_{1}s}\mathrm{d}a+\int_{|a|>1}\frac{\mathrm{d}a}{|a|^{p_{1}(\alpha+s)-N+1}}$
<<$\infty$.< ∞ .
Hence $f\in W_{s}^{\vec{p}}$.
Next we show that $f\not\in W_{s}^{\vec{p}}$ whenever $p_{1}(\alpha+s)\leq N$.
When $a>2$, $a^{2}<|x|^{2}<5a^{2}/4$ and $x_{j}>0$, we have $21+|x|^{2}<3a^{2}/$ 2 and $1+|x+ae_{j}|^{2}=1+|x|^{2}+2ax_{j}+a^{2}>2a^{2}$. Hence
$\int_{\mathbb{R}^{N}}|f(x)-f(x+ae_{j})|^{p_{1}}\mathrm{d}x$ $\geq\int_{\begin{subarray}{c}a^{2}<|x|^{2}<5a^{2}/4\\ x_{j}>0\end{subarray}}|f(x)-f(x+ae_{j})|^{p_{1}}\mathrm{d}x$
$\gtrsim\int_{\begin{subarray}{c}a^{2}<|x|^{2}<5a^{2}/4\\ x_{j}>0\end{subarray}}\frac{\mathrm{d}x}{a^{p_{1}\alpha}}$
$\approx a^{N-p_{1}\alpha}.$
Therefore,
$\int_{\mathbb{R}}\int_{\mathbb{R}^{N}}\frac{|f(x)-f(x+ae_{j})|^{p_{1}}}{|a|^{1+p_{1}s}}\mathrm{d}x\,\mathrm{d}a$ $\gtrsim\int_{a\geq 2}\frac{\mathrm{d}a}{a^{p_{1}(\alpha+s)-N+1}}=\infty.$
(ii)  Now assume that $p_{0}>p_{1}$. Take some $\varepsilon>0$ such that $\varepsilon<\min\{s,N/p_{1}-N/p_{0}\}$. Set $\alpha=N/p_{1}-\varepsilon$. We have $p_{1}\alpha<N<p_{0}\alpha$ and $p_{1}(\alpha+s)=N+p_{1}(s-\varepsilon)>N$. Hence $f\in W_{s}^{\vec{p}}\setminus W_{s}^{p_{1}}$.
(iii)  When $\tilde{s}<N/p_{1}-N/p_{0}$, we have $p_{0}>p_{1}$. Take some constant $\alpha$ such that
$\max\Big{\{}\frac{N}{p_{0}},\frac{N}{p_{1}}-s\Big{\}}<\alpha<\frac{N}{p_{1}}-\tilde{s}.$
Then $p_{0}\alpha>N>p_{1}\alpha$, $p_{1}(\alpha+s)>N$ and $p_{1}(\alpha+\tilde{s})<N$. Hence $f\in W_{s}^{\vec{p}}\setminus W_{\tilde{s}}^{\vec{p}}$. ∎
Proof of Theorem 3.1.
For the case $p_{0}\!\leq p_{1}$, the conclusion follows from Lemma 3.4 and the embedding theorem for classical fractional Sobolev spaces.
Now we assume that $p_{0}>p_{1}$. It suffices to show that if $f\in W_{s}^{\vec{p}}$, then $f\in L^{q}$ with $q=Np_{1}/(N-sp_{1})$.
We adopt the notation used in the proof of Lemma 3.4. We rewrite (3.11) as
$g$ $=((\nabla_{x}\Phi)E)*\nabla_{x}g+(\Phi\nabla_{x}E)*\nabla_{x}g+((\partial_{t}\Phi)E)*h_{0}+(\Phi\partial_{t}E)*h_{0}$
$\qquad+((\partial_{t}\Phi)E)*h_{1}+(\Phi\partial_{t}E)*h_{1}-2(\nabla\Phi\cdot\nabla E)*g-((\Delta\Phi)E)*g.$ (3.17)
Checking the proof of [16, Proposition 4.47], we find that
$\Big{\|}\Big{(}((\nabla_{x}\Phi)E)*\nabla_{x}g+(\Phi\nabla_{x}E)*\nabla_{x}g+((\partial_{t}\Phi)E)*h_{0}+(\Phi\partial_{t}E)*h_{0}\Big{)}(0,\cdot)\Big{\|}_{L^{q}}\lesssim\|f\|_{W_{s}^{\vec{p}}}.$ (3.18)
In fact, the only property of $g$ used in the proof of [16, Proposition 4.47] is that $t^{\alpha}\nabla g(t,x)$ lies in $L^{p_{1}}$. In our case, $\partial_{t}g$ is replaced by $h_{0}$. By Lemma 3.3, $t^{\alpha}h_{0}(t,x)\in L^{p_{1}}$. So (3.18) is true.
It remains to consider the last four terms in (3.17). Since $\Phi$ is compactly supported, we have
$|(\partial_{t}\Phi)E(t,x)|\lesssim\frac{1}{(|t|+|x|)^{N}},$
and
$|\Phi\partial_{t}E(t,x)|\lesssim\frac{1}{(|t|+|x|)^{N}}.$
Observe that $h_{1}(t,x)=0$ for $t<1$ or $t>A$. We have
$\Big{|}\Big{(}((\partial_{t}\Phi)E)*h_{1}+(\Phi\partial_{t}E)*h_{1}\Big{)}(0,x)\Big{|}\lesssim\int_{1}^{A}\int_{[-A,A]^{N}}\frac{|h_{1}(t,x-y)|}{(t+|y|)^{N}}\mathrm{d}y\,\mathrm{d}t.$
Take some $r>1$ such that $11/p_{0}+1/r=1/q+$ 1. Applying Young’s inequality, we deduce from (3.5) that
$\Big{\|}\Big{(}((\partial_{t}\Phi)E)*h_{1}+(\Phi\partial_{t}E)*h_{1}\Big{)}(0,\cdot)\Big{\|}_{L^{q}}\lesssim\int_{1}^{A}\|h_{1}(t,\cdot)\|_{L^{p_{0}}}\Big{\|}\frac{1}{(t+|\cdot|)^{N}}\Big{\|}_{L^{r}}\mathrm{d}t\lesssim\|f\|_{L^{p_{0}}}.$
For the last two terms in (3.17), we show in (S3) of the proof of Lemma 3.4 that
$\Big{\|}\Big{(}|(\nabla\Phi\cdot\nabla E)*g|+|((\Delta\Phi)E)*g|\Big{)}(0,\cdot)\Big{\|}_{L^{q}}\lesssim\|f\|_{L^{p_{0}}}.$
Combining these facts we derive the desired conclusion. ∎
3.2. Density of Compactly Supported Infinitely Differentiable Functions
In this subsection, we show that compactly supported infinitely differentiable functions are dense in fractional nonuniform Sobolev spaces for certain indices.
First, we show that smooth functions are dense in nonuniform fractional Sobolev spaces.
Lemma 3.7.
For any $10<s<$ 1 and $\vec{p}=(p_{0},p_{1})$ with $1\leq p_{0},p_{1}<\infty$, $C^{\infty}\cap W_{s}^{\vec{p}}(\mathbb{R}^{N})$ is dense in $W_{s}^{\vec{p}}(\mathbb{R}^{N})$.
Proof.
As in the proof of Lemma 2.1, take some $\varphi\in C_{c}^{\infty}$ such that $\varphi$ is nonnegative and $\|\varphi\|_{L^{1}}=1$. Fix some $f\in W_{s}^{\vec{p}}$ and Set
$g_{\lambda}(x)=\int_{\mathbb{R}^{N}}f(y)\frac{1}{\lambda^{N}}\varphi(\frac{x-y}{\lambda})\mathrm{d}y,\qquad\lambda>0.$
We have $g_{\lambda}\in C^{\infty}$. Moreover, since
$g_{\lambda}(x)=\int_{\mathbb{R}^{N}}f(x-\lambda y)\varphi(y)\mathrm{d}y,$
we derive from Minkowski’s inequality and the continuity of translation operators in $L^{p_{0}}$ that
$\lim_{\lambda\rightarrow 0}\|g_{\lambda}-f\|_{L^{p_{0}}}=0$ (3.19)
and
$[g_{\lambda}]_{W_{s}^{p_{1}}}$ $=\bigg{\|}\frac{g_{\lambda}(x)-g_{\lambda}(z)}{|x-z|^{N/p_{1}+s}}\bigg{\|}_{L^{p_{1}}(\mathbb{R}^{N}\times\mathbb{R}^{N})}\leq\int_{\mathbb{R}^{N}}[f]_{W_{s}^{p_{1}}}\varphi(y)\mathrm{d}y=[f]_{W_{s}^{p_{1}}},\quad\forall\lambda>0.$
On the other hand, we see from (3.19) that there is a subsequence $\{g_{\lambda_{l}}:\,l\geq 1\}$ which is convergent to $f$ almost everywhere. It follows from Fatou’s lemma that
$[f]_{W_{s}^{p_{1}}}\leq\liminf_{l\rightarrow\infty}[g_{\lambda_{l}}]_{W_{s}^{p_{1}}}\leq\limsup_{l\rightarrow\infty}[g_{\lambda_{l}}]_{W_{s}^{p_{1}}}\leq[f]_{W_{s}^{p_{1}}}.$
Hence
$\lim_{l\rightarrow\infty}[g_{\lambda_{l}}]_{W_{s}^{p_{1}}}=[f]_{W_{s}^{p_{1}}}.$
That is, $\{(g_{\lambda_{l}}(x)-g_{\lambda_{l}}(z))/|x-z|^{N/p_{1}+s}:\,l\geq 1\}$ converges to $(f(x)-f(z))/|x-z|^{N/p_{1}+s}$ both in $L^{p_{1}}$ norm and almost everywhere on $\mathbb{R}^{N}\times\mathbb{R}^{N}$. It follows from Riesz’s theorem that
$\lim_{l\rightarrow\infty}[g_{\lambda_{l}}-f]_{W_{s}^{\vec{p}}}=0.$
Together with (3.19), we get the conclusion as desired. ∎
As in the classical case, to prove the density of compactly supported smooth functions, we first approximate a function in fractional nonuniform Sobolev spaces by its truncation, then approximate the truncation by its regularization. However, since $p_{0}$ needs not to be identical to $p_{1}$, the technique details are quite different.
Theorem 3.8.
Suppose that $10<s<$ 1, $\vec{p}=(p_{0},p_{1})\in[1,\infty)^{2}$ and $s/N\geq 1/p_{1}-1/p_{0}$. Then $C_{c}^{\infty}(\mathbb{R}^{N})$ is dense in $W_{s}^{\vec{p}}(\mathbb{R}^{N})$.
Proof.
Take some nonnegative function $\varphi\in C_{c}^{\infty}$ such that $\varphi(x)=1$ when $|x|<1$ and $\varphi(x)=0$ when $|x|>2$.
Fix some $f\in W_{s}^{\vec{p}}$. First, we show that $\varphi(\cdot/n)f$ tends to $f$ in $W_{s}^{\vec{p}}$ as $n$ tends to the infinity. Since $\varphi(\cdot/n)f$ tends to $f$ in $L^{p_{0}}$, thanks to Lebesgue’s dominated convergence theorem, it suffices to show that $\lim_{n\rightarrow\infty}[\varphi(\cdot/n)f-f]_{W_{s}^{p_{1}}}=0$.
If $p_{0}\leq p_{1}$, by Lemma 3.4, we have $W_{s}^{\vec{p}}\hookrightarrow W_{s}^{p_{1}}$. Now we see from the density result in the classical fractional Sobolev spaces that $\lim_{n\rightarrow\infty}[\varphi(\cdot/n)f-f]_{W_{s}^{p_{1}}}=0$.
It remains to consider the case $p_{0}>p_{1}$. Observe that
$\Big{(}\varphi\big{(}\frac{x}{n}\big{)}-1\Big{)}f(x)-\Big{(}\varphi\big{(}\frac{y}{n}\big{)}-1\Big{)}f(y)=\Big{(}\varphi\big{(}\frac{x}{n}\big{)}-1\Big{)}(f(x)-f(y))+\Big{(}\varphi\big{(}\frac{x}{n}\big{)}-\varphi\big{(}\frac{y}{n}\big{)}\Big{)}f(y).$
We have
$[\varphi\big{(}\frac{\cdot}{n}\big{)}f-f]_{W_{s}^{p_{1}}}^{p_{1}}$ $\lesssim\iint_{\mathbb{R}^{2N}}\frac{|(\varphi(x/n)-1)(f(x)-f(y))|^{p_{1}}}{|x-y|^{N+sp_{1}}}\mathrm{d}x\,\mathrm{d}y$
$\qquad+\iint_{\mathbb{R}^{2N}}\frac{|(\varphi(x/n)-\varphi(y/n))f(y)|^{p_{1}}}{|x-y|^{N+sp_{1}}}\mathrm{d}x\,\mathrm{d}y.$
Since $f\in W_{s}^{\vec{p}}$, we see from Lebesgue’s dominated convergence theorem that
$\lim_{n\rightarrow\infty}\iint_{\mathbb{R}^{2N}}\frac{|(\varphi(x/n)-1)(f(x)-f(y))|^{p_{1}}}{|x-y|^{N+sp_{1}}}\mathrm{d}x\,\mathrm{d}y=0.$
Now we only need to show that
$\lim_{n\rightarrow\infty}\iint_{\mathbb{R}^{2N}}\frac{|(\varphi(x/n)-\varphi(y/n))f(y)|^{p_{1}}}{|x-y|^{N+sp_{1}}}\mathrm{d}x\,\mathrm{d}y=0.$ (3.20)
We split the integral into two parts,
$\iint_{\mathbb{R}^{2N}}\frac{|(\varphi(x/n)-\varphi(y/n))f(y)|^{p_{1}}}{|x-y|^{N+sp_{1}}}\mathrm{d}x\,\mathrm{d}y$
$=\Big{(}\iint_{|y|\leq 4n}+\iint_{|y|>4n}\Big{)}\frac{|(\varphi(x/n)-\varphi(y/n))f(y)|^{p_{1}}}{|x-y|^{N+sp_{1}}}\mathrm{d}x\,\mathrm{d}y$
$=I+II.$
We have
$I$ $=\Big{(}\iint_{\begin{subarray}{c}|y|\leq 4n\\ |x-y|\leq 8n\end{subarray}}+\iint_{\begin{subarray}{c}|y|\leq 4n\\ |x-y|>8n\end{subarray}}\Big{)}\frac{|(\varphi(x/n)-\varphi(y/n))f(y)|^{p_{1}}}{|x-y|^{N+sp_{1}}}\mathrm{d}x\,\mathrm{d}y$
$=I_{1}+I_{2},$
where
$I_{1}$ $\leq\frac{\|\nabla\varphi\|_{L^{\infty}}^{p_{1}}}{n^{p_{1}}}\iint_{\begin{subarray}{c}|y|\leq 4n\\ |x-y|\leq 8n\end{subarray}}\frac{|x-y|^{p_{1}}|f(y)|^{p_{1}}}{|x-y|^{N+sp_{1}}}\mathrm{d}x\,\mathrm{d}y\lesssim\frac{1}{n^{sp_{1}}}\int_{|y|\leq 4n}|f(y)|^{p_{1}}\mathrm{d}y,$
and
$I_{2}$ $=\iint_{\begin{subarray}{c}|y|\leq 4n\\ |x-y|>8n\end{subarray}}\frac{|\varphi(y/n)f(y)|^{p_{1}}}{|x-y|^{N+sp_{1}}}\mathrm{d}x\,\mathrm{d}y\lesssim\frac{1}{n^{sp_{1}}}\int_{|y|\leq 4n}|f(y)|^{p_{1}}\mathrm{d}y.$
Hence
$I\lesssim\frac{1}{n^{sp_{1}}}\int_{|y|\leq 4n}|f(y)|^{p_{1}}\mathrm{d}y.$
Denote $r=p_{0}/p_{1}$. For any $k\geq 1$, there is some $n_{0}$ such that
$\|f\cdot 1_{\{|y|\leq n_{0}\}}\|_{L^{p_{0}}}\geq\Big{(}1-\frac{1}{k}\Big{)}\|f\|_{L^{p_{0}}}.$
Applying Hölder’s inequality, we get
$I$ $\lesssim\frac{1}{n^{sp_{1}}}\Big{(}\int_{|y|\leq n_{0}}|f(y)|^{p_{1}}\mathrm{d}y+\int_{n_{0}<|y|\leq 4n}|f(y)|^{p_{1}}\mathrm{d}y\Big{)}$
$\lesssim\frac{1}{n^{sp_{1}}}\Big{(}n_{0}^{N/r^{\prime}}\||f|^{p_{1}}\|_{L^{r}}+n^{N/r^{\prime}}\||f|^{p_{1}}\cdot 1_{\{|y|>n_{0}\}}\|_{L^{r}}\Big{)}$
$\leq\frac{n_{0}^{N/r^{\prime}}}{n^{sp_{1}}}\|f\|_{L^{p_{0}}}^{p_{1}}+\frac{1}{k^{p_{1}}}\cdot\frac{1}{n^{Np_{1}(s/N-1/p_{1}+1/p_{0})}}\|f\|_{L^{p_{0}}}^{p_{1}}.$
Since $s/N\geq 1/p_{1}-1/p_{0}$, letting $n\rightarrow\infty$ and $k\rightarrow\infty$ successively, we get
$\lim_{n\rightarrow\infty}I=0.$
Next we deal with $II$. Recall that $\varphi(x)=0$ when $|x|>2$. We have
$II$ $=\iint_{|y|>4n}\frac{|(\varphi(x/n)-\varphi(y/n))f(y)|^{p_{1}}}{|x-y|^{N+sp_{1}}}\mathrm{d}x\mathrm{d}y$
$=\int_{|x|\leq 2n}\int_{|y|>4n}\frac{|\varphi(x/n)f(y)|^{p_{1}}}{|x-y|^{N+sp_{1}}}\mathrm{d}x\mathrm{d}y$
$\lesssim\int_{|x|\leq 2n}|\varphi(\frac{x}{n})|^{p_{1}}\mathrm{d}x\int_{|y|>4n}\frac{|f(y)|^{p_{1}}}{|y|^{N+sp_{1}}}\mathrm{d}y$
$\lesssim n^{N}\cdot\|f\cdot 1_{\{|y|\geq 4n\}}\|_{L^{p_{0}}}^{p_{1}}\cdot\frac{1}{n^{sp_{1}+N/r}}$
$=\frac{1}{n^{Np_{1}(s/N-1/p_{1}+1/p_{0})}}\|f\cdot 1_{\{|y|\geq 4n\}}\|_{L^{p_{0}}}^{p_{1}}.$
Hence $\lim_{n\rightarrow\infty}II=0$. Therefore, (3.20) is true.
The above arguments show that the sequence $\{\varphi(\cdot/n)f\>\,n\geq 1\}$ is a subset of $W_{s}^{\vec{p}}$ and is convergent to $f$ in $W_{s}^{\vec{p}}$.
To finish the proof, it suffices to show that for any compactly supported function $f\in W_{s}^{\vec{p}}$, $f$ can be approximated by functions in $C_{c}^{\infty}$, which can be achieved with the same arguments as in the proof of Lemma 3.7. ∎
4. Convergence of Schrödinger Operators
In this section, we study the convergence of Schrödinger operators.
Take some function $\varphi$ such that $\hat{\varphi}\in C_{c}^{\infty}$, $0\leq\hat{\varphi}(\omega)\leq 1$, $\hat{\varphi}(\omega)=1$ for $|\omega|<1$ and $\hat{\varphi}(\omega)=0$ for $|\omega|>2$. Set $\hat{f}_{1}=\hat{\varphi}\cdot\hat{f}$ and $f_{2}=f-f_{1}$.
Note that $(f*\varphi)\,\hat{}=\hat{\varphi}\hat{f}$ (for a proof, see [28, Theorem 7.19]). Hence $f_{1}=f*\varphi$ and $f_{2}=f-f*\varphi$. Now we rewrite $e^{it(-\Delta)^{a/2}}$ as
$e^{it(-\Delta)^{a/2}}f=e^{it(-\Delta)^{a/2}}f_{1}+e^{it(-\Delta)^{a/2}}f_{2}.$ (4.1)
If $f\in H^{s}$, then $\hat{f}$ is locally integrable. Hence the convergence for
$\lim_{t\rightarrow 0}e^{it(-\Delta)^{a/2}}f_{1}(x)=\lim_{t\rightarrow 0}\frac{1}{(2\pi)^{N}}\int_{\mathbb{R}^{N}}e^{i(x\cdot\omega+t|\omega|^{a})}\hat{f}_{1}(\omega)\mathrm{d}\omega$
is obvious.
However, for $f\in W_{s}^{\vec{p}}$ with $p_{0}>2$, we do not know whether $\hat{f}$ is locally integrable. So we have to deal with the term $e^{it(-\Delta)^{a/2}}f_{1}$ with new method.
For the case $p_{\lceil s\rceil}=2$, we show that if $e^{it(-\Delta)^{a/2}}f$ is convergent as $t$ tends to zero for all functions in $H^{s}$ for some $s>0$, then the same is true for all functions in $W_{s}^{\vec{p}}$ with the same index $s$.
Theorem 4.1.
Let $s>0$, $a>1$, $\vec{p}=(p_{0},\ldots,p_{\lceil s\rceil})$ with $p_{\lceil s\rceil}=2$ and $1\leq p_{l}<\infty$ for all $10\leq l\leq\lceil s\rceil-$ 1. Suppose that for all functions $f\in H^{s}(\mathbb{R}^{N})$,
$\lim_{t\rightarrow 0}e^{it(-\Delta)^{a/2}}f(x)=f(x),\quad a.e.$ (4.2)
We have:
  1. (i)
    If $20<s<N/$ 2, then (4.2) is true for all functions $f\in W_{s}^{\vec{p}}$.
  2. (ii)
    If $s\geq N/2$ and $a=2$, then (4.2) is also true if we interpret the operator $e^{-it\Delta}$ as
    $e^{-it\Delta}f(x)=\lim_{\varepsilon\rightarrow 0}\frac{1}{(2\pi)^{N}}\int_{\mathbb{R}^{N}}e^{i(x\cdot\omega+t\omega^{2})}\Big{(}1-\varphi_{0}(\frac{\omega}{\varepsilon})\Big{)}\hat{f}(\omega)\mathrm{d}\omega,$ (4.3)
    where $\varphi_{0}\in C_{c}^{\infty}(\mathbb{R})$ satisfies that $\varphi_{0}(x)=1$ for $|x|<1$ and $\varphi_{0}(x)=0$ for $|x|>2$, the limit exists almost everywhere, and the limit is independent of the choice of $\varphi_{0}$ with the aforementioned properties.
For the case $21<p_{\lceil s\rceil}<$ 2 and $a=2$, we have similar results.
Theorem 4.2.
Let $s>0$, $\vec{p}=(p_{0},\ldots,p_{\lceil s\rceil})$ with $21<p_{\lceil s\rceil}<$ 2 and $1\leq p_{l}<\infty$ for all $10\leq l\leq\lceil s\rceil-$ 1. Let
$\beta_{s}=\begin{cases}s,&s\in\mathbb{Z},\\ \lfloor s\rfloor+\nu_{s}(p_{\lceil s\rceil}-1),&s\not\in\mathbb{Z}.\end{cases}$ (4.4)
We have:
  1. (i)
    If
    $\frac{1}{p_{\lceil s\rceil}}-\frac{N}{2(N+1)}<\frac{\beta_{s}}{N}<\frac{1}{p_{\lceil s\rceil}},$ (4.5)
    then for any $f\in W_{s}^{\vec{p}}$,
    $\lim_{t\rightarrow 0}e^{-it\Delta}f(x)=f(x),\quad a.e.$ (4.6)
  2. (ii)
    If $\beta_{s}\geq N/p_{\lceil s\rceil}$, then (4.6) is also true if we interpret the operator $e^{-it\Delta}$ as in (4.3).
Remark 4.3.
If $10<s<$ 1, $p_{0}>N/\alpha>p_{1}=2$ and $2(\alpha+s)>N$, then the function $f$ defined in Example 3.6 satisfies that $f\in W_{s}^{(p_{0},2)}\setminus W_{s}^{(2,2)}$. That is, $W_{s}^{(p_{0},2)}$ is not contained in $W_{s}^{(2,2)}=H^{s}$.
For the case $p_{0}>2$, the Fourier transform of functions in $W_{s}^{\vec{p}}$ might be distributions. In this case, we show that $\hat{f}$ is the distributional limit of a sequence of locally integrable functions. Moreover, we prove that $\hat{f}$ is locally integrable if $\beta_{s}/N<1/p_{\lceil s\rceil}$.
Lemma 4.4.
Suppose that $s>0$ and $\vec{p}=(p_{0},\ldots,p_{\lceil s\rceil})$ with $1<p_{\lceil s\rceil}\leq 2$ and $1\leq p_{l}<\infty$ for $10\leq l\leq\lceil s\rceil-$ 1. Let $\beta_{s}$ be defined by (4.4). For any $f\in W_{s}^{\vec{p}}$, $\hat{f}$ coincides with a function $\lambda$ in the domain $\mathbb{R}^{N}\setminus\{0\}$ such that $|x|^{\beta_{s}}\lambda(x)\in L^{p^{\prime}_{\lceil s\rceil}}$ and $\hat{f}=\lim_{\varepsilon\rightarrow 0}(1-\varphi_{0}(x/\varepsilon))\lambda(x)$ in the distributional sense, where $\varphi_{0}\in C^{\infty}(\mathbb{R}^{N})$ satisfies that $\varphi_{0}(x)=1$ for $|x|<1$ and $\varphi_{0}(x)=0$ for $|x|>2$.
Consequently, if $\beta_{s}p_{\lceil s\rceil}<N$, then $\lambda$ is locally integrable and $\hat{f}=\lambda$.
Proof.
First, we show that there is some function $\lambda$ such that $|x|^{\beta_{s}}\lambda(x)\in L^{p^{\prime}_{\lceil s\rceil}}$ and
$\langle\hat{f},\varphi\rangle=\int_{\mathbb{R}^{N}}\lambda(x)\varphi(x)\mathrm{d}x,\qquad\forall\varphi\in C_{c}^{\infty}(\mathbb{R}^{N}\setminus\{0\}).$ (4.7)
There are three cases.
(A1)  $10<s<$ 1.
In this case, $\vec{p}=(p_{0},p_{1})$. We see from Proposition 3.2 that
$\iint_{\mathbb{R}^{N}\times\mathbb{R}^{N}}\frac{|f(x)-f(y)|^{p_{1}}}{|x-y|^{N+p_{1}s}}\mathrm{d}x\,\mathrm{d}y<\infty$
is equivalent to
$\iint_{\mathbb{R}^{N}\times\mathbb{R}}\frac{|f(x)-f(x+ae_{j})|^{p_{1}}}{|a|^{1+p_{1}s}}\mathrm{d}x\,\mathrm{d}a<\infty,\qquad\forall 1\leq j\leq N.$ (4.8)
Applying Fubini’s theorem, we get that for almost all real numbers $a$, $f-f(\cdot+ae_{j})\in L^{p_{1}}$. It is easy to see that $f(\cdot-ae_{j})\,\hat{}=\phi_{j}(a\cdot)\hat{f}$ in the distributional sense, where
$\phi_{j}(x)=e^{-ix_{j}}.$
Hence
$(f-f(\cdot+ae_{j}))\,\hat{}=(1-\phi_{j}(a\cdot))\hat{f}.$
Since $f-f(\cdot+ae_{j})\in L^{p_{1}}$ for almost all $a>0$, by the Hausdorff-Young inequality, there exist functions $g_{a}\in L^{p^{\prime}_{1}}$ such that
$(f-f(\cdot+ae_{j}))\,\hat{}=g_{a}.$
Hence for any $\varphi\in C_{c}^{\infty}(B(0,1/a)\setminus\{0\})$,
$\langle(1-\phi_{j}(a\cdot))\hat{f},\varphi\rangle$ $=\int_{\mathbb{R}^{N}}g_{a}(x)\varphi(x)\mathrm{d}x.$
Therefore,
$\langle|1-\phi_{j}(a\cdot)|^{2}\hat{f},\varphi\rangle$ $=\langle(1-\phi_{j}(a\cdot))\hat{f},(1-\phi_{j}(a\cdot))^{*}\varphi\rangle$ $=\int_{\mathbb{R}^{N}}(1-\phi_{j}(ax))^{*}g_{a}(x)\varphi(x)\mathrm{d}x,$
where $z^{*}$ denotes the conjugate of a complex number $z$. It follows that
$\Big{\langle}\sum_{j=1}^{N}|1-\phi_{j}(a\cdot)|^{2}\hat{f},\varphi\Big{\rangle}$ $=\int_{\mathbb{R}^{N}}\sum_{j=1}^{N}(1-\phi_{j}(ax))^{*}g_{a}(x)\varphi(x)\mathrm{d}x,$
Set
$\lambda_{a}(x)=\frac{\sum_{j=1}^{N}(1-\phi_{j}(ax))^{*}g_{a}(x)}{\sum_{j=1}^{N}|1-\phi_{j}(ax)|^{2}}.$ (4.9)
Note that $\sum_{j=1}^{N}|1-\phi_{j}(a\cdot)|^{2}$ has no zero in the area $0<|x|<1/a$. Hence
$\frac{\varphi}{\sum_{j=1}^{N}|1-\phi_{j}(a\cdot)|^{2}}\in C_{c}^{\infty}(\mathbb{R}^{N}\setminus\{0\}).$
It follows that
$\Big{\langle}\sum_{j=1}^{N}|1-\phi_{j}(a\cdot)|^{2}\hat{f},\frac{\varphi}{\sum_{j=1}^{N}|1-\phi_{j}(a\cdot)|^{2}}\Big{\rangle}=\int_{\mathbb{R}^{N}}\lambda_{a}(x)\varphi(x)\mathrm{d}x.$
On the other hand,
$\Big{\langle}\sum_{j=1}^{N}|1-\phi_{j}(a\cdot)|^{2}\hat{f},\frac{\varphi}{\sum_{j=1}^{N}|1-\phi_{j}(a\cdot)|^{2}}\Big{\rangle}$ $=\langle\hat{f},\varphi\rangle.$
Hence
$\langle\hat{f},\varphi\rangle=\int_{\mathbb{R}^{N}}\lambda_{a}(x)\varphi(x)\mathrm{d}x,\qquad\forall\varphi\in C_{c}^{\infty}(B(0,1/a)\setminus\{0\}).$ (4.10)
That is, $\hat{f}$ coincides with a function $\lambda_{a}$ in the domain $B(0,1/a)\setminus\{0\}$. By the uniqueness, we have
$\lambda_{a}(x)=\lambda_{a^{\prime}}(x),\qquad a^{\prime}<a,0<|x|<\frac{1}{a}.$
Take a sequence $\{a_{n}:\,n\geq 1\}$ such that $a_{n}\rightarrow 0$. Then the limit
$\lambda(x):=\lim_{n\rightarrow\infty}\lambda_{a_{n}}(x)$
exists for almost all $x\neq 0$. Now we see from (4.10) that (4.7) is true.
Since $f-f(\cdot+ae_{j})\in L^{p_{1}}$, we see from the above equation that for all $a\in\mathbb{R}$,
$\Big{(}f-f(\cdot+ae_{j})\Big{)}\,\hat{}\,(x)=\lambda(x)(1-\phi_{j}(ax)),\qquad\forall\varepsilon<|x|<A.$
Applying the Hausdorff Young inequality, we get
$\iint_{\mathbb{R}^{N}\times\mathbb{R}}\frac{|f(x)-f(x+ae_{j})|^{p_{1}}}{|a|^{1+p_{1}s}}\mathrm{d}x\,\mathrm{d}a$
$\geq\iint_{\mathbb{R}^{N}\times\mathbb{R}}\frac{|(f-f(\cdot+ae_{j}))\,\hat{}(x)|^{p^{\prime}_{1}}}{|a|^{1+p_{1}s}}\mathrm{d}x\,\mathrm{d}a$
$=\lim_{\begin{subarray}{c}\varepsilon\rightarrow 0\\ A\rightarrow\infty\end{subarray}}\int_{\varepsilon<|x|<A}\int_{\mathbb{R}}\frac{|(f-f(\cdot+ae_{j}))\,\hat{}(x)|^{p^{\prime}_{1}}}{|a|^{1+p_{1}s}}\mathrm{d}a\,\mathrm{d}x$ (4.13)
$=\lim_{\begin{subarray}{c}\varepsilon\rightarrow 0\\ A\rightarrow\infty\end{subarray}}\int_{\varepsilon<|x|<A}\int_{\mathbb{R}}\frac{|\lambda(x)(1-\phi_{j}(ax))|^{p^{\prime}_{1}}}{|a|^{1+p_{1}s}}\mathrm{d}a\,\mathrm{d}x$ (4.16)
$=\int_{\mathbb{R}^{N}}\int_{\mathbb{R}}\frac{|\lambda(x)|^{p^{\prime}_{1}}|2\sin(ax_{j}/2)|^{p^{\prime}_{1}}}{|a|^{1+p_{1}s}}\mathrm{d}a\,\mathrm{d}x$
$=\int_{\mathbb{R}^{N}}|x_{j}|^{p_{1}s}|\lambda(x)|^{p^{\prime}_{1}}\int_{\mathbb{R}}\frac{|2\sin(a/2)|^{p^{\prime}_{1}}}{|a|^{1+p_{1}s}}\mathrm{d}a\,\mathrm{d}x$
$=C_{s,j}\int_{\mathbb{R}^{N}}|x_{j}|^{p_{1}s}|\lambda(x)|^{p^{\prime}_{1}}\mathrm{d}x.$ (4.17)
Hence
$\int_{\mathbb{R}^{N}}|x_{j}|^{p_{1}s}|\lambda(x)|^{p^{\prime}_{1}}\mathrm{d}x\leq\frac{1}{C_{s,j}}\iint_{\mathbb{R}^{N}\times\mathbb{R}}\frac{|f(x)-f(x+ae_{j})|^{p_{1}}}{|a|^{1+p_{1}s}}\mathrm{d}x\,\mathrm{d}a<\infty.$ (4.18)
That is, $|x_{j}|^{s(p_{1}-1)}\lambda(x)\in L^{p^{\prime}_{1}}$ for all $1\leq j\leq N$. Consequently, $|x|^{s(p_{1}-1)}\lambda(x)\in L^{p^{\prime}_{1}}$.
(A2)  $s=k\geq 1$ is an integer.
For all multi-index $\alpha$ with $|\alpha|=k$, let $h_{\alpha}$ be the Fourier transform of $D^{\alpha}f$ and $\psi_{\alpha}(x)=(ix)^{\alpha}$. Then $h_{\alpha}=\psi_{\alpha}\hat{f}$ is a function in $L^{p^{\prime}_{k}}$. Consequently,
$|\omega|^{s}|\hat{f}(\omega)|\approx\sum_{|\alpha|=k}|h_{\alpha}(\omega)|\in L^{p^{\prime}_{k}}.$
For any $\varphi\in C_{c}^{\infty}(\mathbb{R}^{N})$ with $\varphi(x)=0$ in a neighbourhood of 00, since $h_{\alpha}=\psi_{\alpha}\hat{f}$ is a function in $L^{p^{\prime}_{k}}$, we have
$\langle\psi_{\alpha}\hat{f},\varphi\rangle=\int_{\mathbb{R}^{N}}h_{\alpha}(x)\varphi(x)\mathrm{d}x.$
Hence
$\langle\psi_{\alpha}^{2}\hat{f},\varphi\rangle=\langle\psi_{\alpha}\hat{f},\psi_{\alpha}\varphi\rangle=\int_{\mathbb{R}^{N}}h_{\alpha}(x)\psi_{\alpha}(x)\varphi(x)\mathrm{d}x.$
Therefore,
$\Big{\langle}\sum_{|\alpha|=k}\psi_{\alpha}^{2}\hat{f},\varphi\Big{\rangle}=\int_{\mathbb{R}^{N}}\sum_{|\alpha|=k}h_{\alpha}(x)\psi_{\alpha}(x)\varphi(x)\mathrm{d}x.$
Note that $\sum_{|\alpha|=k}\psi_{\alpha}^{2}(x)$ has no zero other than $x=0$. Substituting $\varphi/\sum_{|\alpha|=k}\psi_{\alpha}^{2}$ for $\varphi$ in the above equation, we get
$\langle\hat{f},\varphi\rangle=\Big{\langle}\sum_{|\alpha|=k}\psi_{\alpha}^{2}\hat{f},\frac{\varphi}{\sum_{|\alpha|=k}\psi_{\alpha}^{2}}\Big{\rangle}=\int_{\mathbb{R}^{N}}\frac{\sum_{|\alpha|=k}h_{\alpha}(x)\psi_{\alpha}(x)}{\sum_{|\alpha|=k}\psi_{\alpha}(x)^{2}}\varphi(x)\mathrm{d}x.$
Let
$\lambda(x)=\frac{\sum_{|\alpha|=k}h_{\alpha}(x)\psi_{\alpha}(x)}{\sum_{|\alpha|=k}\psi_{\alpha}(x)^{2}}.$
We get (4.7) and $|x|^{s}|\lambda(x)|\lesssim\sum_{\alpha}|h_{\alpha}(x)|\in L^{p^{\prime}_{\lceil s\rceil}}$.
(A3)  $s>1$ is not an integer.
For any multi-index $\alpha$ with $|\alpha|=\lfloor s\rfloor$, set $\psi_{\alpha}(x)=(ix)^{\alpha}$. Substituting $(p_{\lceil s\rceil-1},p_{\lceil s\rceil})$, $\nu_{s}$ and $D^{\alpha}f$ for $(p_{0},p_{1})$, $s$ and $f$ respectively in (A1), we get $|\omega|^{\nu_{s}(p_{\lceil s\rceil}-1)}(D^{\alpha}f)\,\hat{}(\omega)\in L^{p^{\prime}_{\lceil s\rceil}}$ and there exists a function $\lambda_{\alpha}$ such that $|\omega|^{\nu_{s}(p_{\lceil s\rceil}-1)}\lambda_{\alpha}(\omega)\in L^{p^{\prime}_{\lceil s\rceil}}$ and for any $\varphi\in C_{c}^{\infty}(\mathbb{R}^{N})$ with $\varphi(x)=0$ in a neighbourhood of 00,
$\langle\psi_{\alpha}\hat{f},\varphi\rangle$ $=\langle(D^{\alpha}f)\,\hat{},\varphi\rangle=\int_{\mathbb{R}^{N}}\lambda_{\alpha}(x)\varphi(x)\mathrm{d}x.$
With similar arguments as in the previous case, we get
$\langle\hat{f},\varphi\rangle=\int_{\mathbb{R}^{N}}\frac{\sum_{|\alpha|=\lfloor s\rfloor}\lambda_{\alpha}(x)\psi_{\alpha}(x)}{\sum_{|\alpha|=\lfloor s\rfloor}\psi_{\alpha}(x)^{2}}\varphi(x)\mathrm{d}x.$
Let
$\lambda(x)=\frac{\sum_{|\alpha|=\lfloor s\rfloor}\lambda_{\alpha}(x)\psi_{\alpha}(x)}{\sum_{|\alpha|=\lfloor s\rfloor}\psi_{\alpha}(x)^{2}}.$
We get (4.7) and $|x|^{\lfloor s\rfloor+\nu_{s}(p_{\lceil s\rceil}-1)}\lambda(x)|\lesssim\sum_{\alpha}|x|^{\nu_{s}(p_{\lceil s\rceil}-1)}|\lambda_{\alpha}(x)|\in L^{p^{\prime}_{\lceil s\rceil}}$.
Recall that $C_{c}^{\infty}$ is dense in $\mathscr{S}$. For any $\varphi\in\mathscr{S}$ with $\varphi(x)=0$ in a neighbourhood of 00, we see from (4.7) that
$\langle\hat{f},\varphi\rangle=\int_{\mathbb{R}^{N}}\lambda(x)\varphi(x)\mathrm{d}x.$ (4.19)
Take some $\varphi_{0}\in C^{\infty}(\mathbb{R}^{N})$ such that $\varphi_{0}(x)=1$ for $|x|<1$ and $\varphi_{0}(x)=0$ for $|x|>2$. We conclude that
$\lim_{\varepsilon\rightarrow 0}\varphi_{0}\big{(}\frac{\cdot}{\varepsilon}\big{)}\hat{f}=0,\qquad\mbox{in}\quad\mathscr{S}^{\prime}.$
It suffices to show that for any $\varphi\in\mathscr{S}$,
$\lim_{\varepsilon\rightarrow 0}\langle\varphi_{0}\big{(}\frac{\cdot}{\varepsilon}\big{)}\hat{f},\varphi\rangle=0.$ (4.20)
In fact,
$|\langle\varphi_{0}\big{(}\frac{\cdot}{\varepsilon}\big{)}\hat{f},\varphi\rangle|$ $=|\langle\hat{f},\varphi_{0}\big{(}\frac{\cdot}{\varepsilon}\big{)}\varphi\rangle|$
$=\Big{|}\Big{\langle}f,\Big{(}\varphi_{0}\big{(}\frac{\cdot}{\varepsilon}\big{)}\varphi\Big{)}\,\hat{}\,\Big{\rangle}\Big{|}$
$=\Big{|}\Big{\langle}f,\varepsilon^{N}\hat{\varphi}_{0}(\varepsilon\cdot)*\hat{\varphi}\Big{\rangle}\Big{|}$
$\leq\|f\|_{L^{p_{0}}}\cdot\|\varepsilon^{N}\hat{\varphi}_{0}(\varepsilon\cdot)*\hat{\varphi}\|_{L^{p^{\prime}_{0}}}$
$\leq\|f\|_{L^{p_{0}}}\cdot\|\varepsilon^{N}\hat{\varphi}_{0}(\varepsilon\cdot)\|_{L^{p^{\prime}_{0}}}\cdot\|\hat{\varphi}\|_{L^{1}}$
$\rightarrow 0.$
Hence (4.20) is true. Consequently,
$\hat{f}=\lim_{\varepsilon\rightarrow 0}\Big{(}1-\varphi_{0}\big{(}\frac{\cdot}{\varepsilon}\big{)}\Big{)}\hat{f}\qquad\mbox{in}\quad\mathscr{S}^{\prime}.$
It follows that for any $\varphi\in\mathscr{S}$,
$\langle\hat{f},\varphi\rangle$ $=\lim_{\varepsilon\rightarrow 0}\Big{\langle}\Big{(}1-\varphi_{0}\big{(}\frac{\cdot}{\varepsilon}\big{)}\Big{)}\hat{f},\varphi\Big{\rangle}$
$=\lim_{\varepsilon\rightarrow 0}\Big{\langle}\hat{f},\Big{(}1-\varphi_{0}\big{(}\frac{\cdot}{\varepsilon}\big{)}\Big{)}\varphi\Big{\rangle}$
$=\lim_{\varepsilon\rightarrow 0}\int_{\mathbb{R}^{N}}\lambda(x)\Big{(}1-\varphi_{0}\big{(}\frac{x}{\varepsilon}\big{)}\Big{)}\varphi(x)\mathrm{d}x$
$=\lim_{\varepsilon\rightarrow 0}\int_{|x|>\varepsilon}\lambda(x)\Big{(}1-\varphi_{0}\big{(}\frac{x}{\varepsilon}\big{)}\Big{)}\varphi(x)\mathrm{d}x.$ (4.21)
Hence $\hat{f}=\lim_{\varepsilon\rightarrow 0}(1-\varphi_{0}(x/\varepsilon))\lambda(x)$ in distributional sense.
Recall that $|x|^{\beta_{s}}\lambda(x)\in L^{p^{\prime}_{\lceil s\rceil}}$. If $\beta_{s}p_{\lceil s\rceil}<N$, we see from Hölder’s inequality that $\lambda(x)=|x|^{-\beta_{s}}|x|^{\beta_{s}}\lambda(x)\in L^{1}(B(0,R))$ for any $R>0$. Hence $\lambda(x)$ is locally integrable. This completes the proof. ∎
We are now ready to prove the main results.
Proof of Theorem 4.1.
Fix some $f\in W_{s}^{\vec{p}}$. By Lemma 4.4, there is some $h\in L^{2}$ such that $\hat{f}(\omega)=h(\omega)/|\omega|^{s}$.
Take some function $\varphi$ such that $\hat{\varphi}\in C_{c}^{\infty}$, $0\leq\hat{\varphi}(\omega)\leq 1$, $\hat{\varphi}(\omega)=1$ for $|\omega|<1$ and $\hat{\varphi}(\omega)=0$ for $|\omega|>2$. Set $\hat{f}_{1}=\hat{\varphi}\cdot\hat{f}$ and $f_{2}=f-f_{1}$. Then the decomposition (4.1) is true.
First, we consider the case $s<N/2$. In this case, we have
$\hat{f}_{1}(\omega)=\frac{1}{|\omega|^{s}}1_{\{|\omega|<2\}}(\omega)\cdot\varphi(\omega)h(\omega)\in L^{1}.$
It follows from the dominated convergence theorem that $e^{it(-\Delta)^{a/2}}f_{1}(x)$ tends to $f_{1}(x)$ as $t$ tends to 00.
On the other hand, since $\hat{f}_{2}(\omega)=(1-\hat{\varphi}(\omega))\hat{f}(\omega)=0$ for $|\omega|\leq 1$, we have
$\|\hat{f}_{2}\|_{L^{2}}\leq\||\omega|^{s}\hat{f}_{2}(\omega)\|_{L^{2}}\leq\||\omega|^{s}\hat{f}(\omega)\|_{L^{2}}<\infty.$
Hence $f_{2}\in H^{s}$. By the hypothesis, we get $\lim_{t\rightarrow 0}e^{it(-\Delta)^{a/2}}f_{2}(x)=f_{2}(x)$, a.e.
Next we consider the case $s\geq N/2$ and $a=2$.
Set $\varphi_{0}=\varphi$. Since $01-\varphi_{0}(\omega/\varepsilon)=$ 0 for $|\omega|<\varepsilon$, we rewrite $e^{-it\Delta}f$ as
$e^{-it\Delta}f(x)$ $=\lim_{\varepsilon\rightarrow 0}\frac{1}{(2\pi)^{N}}\int_{|\omega|>\varepsilon}e^{i(x\omega+t\omega^{2})}\Big{(}1-\varphi_{0}\big{(}\frac{\omega}{\varepsilon}\big{)}\Big{)}\hat{f}(\omega)\mathrm{d}\omega.$ (4.22)
Recall that $|\omega|^{s}\hat{f}(\omega)\in L^{2}$. Hence $\hat{f}\cdot 1_{\{|\omega|>\varepsilon\}}\in L^{2}$. Therefore, the integral in the above equation is well defined. Let us show that the limit in (4.22) exists almost everywhere. Observe that
$\int_{|\omega|>\varepsilon}e^{i(x\omega+t\omega^{2})}\Big{(}1-\varphi_{0}\big{(}\frac{\omega}{\varepsilon}\big{)}\Big{)}\hat{f}(\omega)\mathrm{d}\omega$
$=\int_{|\omega|>\varepsilon}e^{i(x\omega+t\omega^{2})}\Big{(}1-\varphi_{0}\big{(}\frac{\omega}{\varepsilon}\big{)}\Big{)}\hat{\varphi}(\omega)\hat{f}(\omega)\mathrm{d}\omega+\int_{|\omega|>1}e^{i(x\omega+t\omega^{2})}\hat{f}_{2}(\omega)\mathrm{d}\omega.$
By (4.21), we have
$e^{-it\Delta}f_{1}(x)$ $=\lim_{\varepsilon\rightarrow 0}\frac{1}{(2\pi)^{N}}\int_{|\omega|>\varepsilon}e^{i(x\omega+t\omega^{2})}\Big{(}1-\varphi_{0}\big{(}\frac{\omega}{\varepsilon}\big{)}\Big{)}\hat{\varphi}(\omega)\hat{f}(\omega)\mathrm{d}\omega$
$=\frac{1}{(2\pi)^{N}}\langle\hat{f},e^{i(x\omega+t\omega^{2})}\hat{\varphi}(\omega)\rangle.$
That is, the limit in (4.22) exists for almost all $x$.
Observe that $e^{i(x\omega+t\omega^{2})}\hat{\varphi}(\omega)$ tends to $e^{ix\omega}\hat{\varphi}(\omega)$ in $\mathscr{S}$ as $t\rightarrow 0$. We get
$\lim_{t\rightarrow 0}e^{-it\Delta}f_{1}(x)$ $=\lim_{t\rightarrow 0}\frac{1}{(2\pi)^{N}}\langle\hat{f},e^{i(x\omega+t\omega^{2})}\hat{\varphi}(\omega)\rangle$
$=\frac{1}{(2\pi)^{N}}\langle\hat{f},e^{ix\omega}\hat{\varphi}(\omega)\rangle$
$=\langle f,\varphi(x-\cdot)\rangle$
$=f*\varphi(x).$
As in the previous case, $f_{2}\in H^{s}$. By the hypothesis, $\lim_{t\rightarrow 0}e^{it(-\Delta)^{a/2}}f_{2}(x)=f_{2}(x)$, a.e. Hence $\lim_{t\rightarrow 0}e^{it(-\Delta)^{a/2}}f(x)=f(x)$, a.e. This completes the proof. ∎
Proof of Theorem 4.2.
Fix some $f\in W_{s}^{\vec{p}}$. By Lemma 4.4, there is some $h\in L^{p^{\prime}_{\lceil s\rceil}}$ such that $\hat{f}(\omega)=h(\omega)/|\omega|^{\beta_{s}}$.
As in the proof of Theorem 4.1, we apply the decomposition (4.1) with $a=2$ for $e^{-it\Delta}f$.
First, we deal with $e^{-it\Delta}f_{2}(x)$. Since $\hat{f}_{2}(\omega)=(1-\hat{\varphi}(\omega))\hat{f}(\omega)=0$ for $|\omega|\leq 1$, we have
$\begin{split}\|\hat{f}_{2}\|_{L^{2}}&=\||\omega|^{-\beta_{s}}\cdot 1_{\{|\omega|>1\}}\cdot|\omega|^{\beta_{s}}\hat{f}(\omega)\|_{L^{2}}\\ &\leq\||\omega|^{-\beta_{s}}\cdot 1_{\{|\omega|>1\}}\|_{L^{r}}\cdot\||\omega|^{\beta_{s}}\hat{f}(\omega)\|_{L^{p^{\prime}_{\lceil s\rceil}}}<\infty,\end{split}$ (4.23)
where $r>1$ satisfying $1/r=1/2-1/p^{\prime}_{\lceil s\rceil}=1/p_{\lceil s\rceil}-1/2<\beta_{s}/N$. More precisely, since
$\frac{1}{p_{\lceil s\rceil}}-\frac{1}{2}+\frac{1}{2(N+1)}=\frac{1}{p_{\lceil s\rceil}}-\frac{N}{2(N+1)}<\frac{\beta_{s}}{N}.$
We have
$\frac{1}{r}<\frac{\beta_{s}-N/(2(N+1))}{N}.$
Consequently, there is some $\tau>N/(2(N+1))$ such that
$\frac{1}{r}<\frac{\beta_{s}-\tau}{N}.$
Applying Hölder’s inequality yields
$\||\omega|^{\tau}\hat{f}(\omega)\|_{L^{2}}\leq\Big{\|}|\omega|^{-(\beta-\tau)}1_{\{|\omega|>1\}}\Big{\|}_{L^{r}}\cdot\Big{\|}|\omega|^{\beta_{s}}\hat{f}(\omega)\Big{\|}_{L^{p^{\prime}_{\lceil s\rceil}}}<\infty.$
Thus $f_{2}\in H^{\tau}$.
Recall that it was proved in [9, 18, 19] that $\lim_{t\rightarrow 0}e^{-it\Delta}f(x)=f(x)$, a.e., for all $f\in H^{\tau}$ with $\tau>N/(2(N+1))$. Hence
$\lim_{t\rightarrow 0}e^{-it\Delta}f_{2}(x)=f_{2}(x),\quad a.e.$
Next we deal with $e^{-it\Delta}f_{1}$. If $\beta_{s}<N/p_{\lceil s\rceil}$, then we have $\hat{f}_{1}\in L^{1}$. It follows from the dominated convergence theorem that $e^{-it\Delta}f_{1}(x)$ tends to $f_{1}(x)$ as $t$ tends to 00.
If $\beta_{s}\geq N/p_{\lceil s\rceil}$, employing the same arguments as in the case $s\geq N/2$ in the proof of Theorem 4.1, we obtain
$\lim_{t\rightarrow 0}e^{-it\Delta}f_{1}(x)$ $=f*\varphi(x).$
Hence $\lim_{t\rightarrow 0}e^{-it\Delta}f(x)=f(x)$, a.e. This completes the proof. ∎